### v2

# Lexical categories of stem-forming roots in Mapudüngun verb forms

Abstract
After developing a computational system for morphological analysis of the Mapuche language, and evaluating it with texts from various authors and styles, it became necessary to verify the linguistic assumptions of the source used as the basis for implementing this tool.
In the present work, the primary focus is on the lexical category classification of Mapudüngun roots recognised as verbal in the source utilised for the development of the morphological analysis system.
The results of this lexical category revision directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, it is hoped that these results will help clarify some uncertainties about lexical categories in the Mapuche language.
This work addresses a preliminary task to identify the valency of true verbal roots, the results of which will be presented in a subsequent work that complements this article.
Resumen
Tras haber desarrollado un sistema computacional para el análisis morfológico de la lengua mapuche y de evaluarlo con textos de diversos autores y estilos, se ha visto la necesidad de verificar los postulados lingüísticos de la fuente utilizada como base para implementar esta herramienta.
En el presente trabajo, el enfoque principal es la clasificación de la categoría léxica de las raíces del mapudüngun reconocidas como verbales en la fuente utilizada para el desarrollo del sistema de análisis morfológico.
Los resultados de esta revisión categorial benefician directamente al analizador computacional, ya que se implementan tan pronto como se verifican. Además, se espera que estos resultados ayuden a aclarar algunas incertidumbres sobre las categorías léxicas en la lengua mapuche.
Este trabajo aborda una tarea preliminar a la identificación de la valencia de las raíces verbales, cuyos resultados se presentarán en un trabajo posterior que complementa este artículo.
Resum
Després d’haver desenvolupat un sistema computacional d’anàlisi morfològica de la llengua mapuche, i d’avaluar-lo amb textos de diversos autors i estils, es va veure la necessitat de verificar els postulats lingüístics de la font utilitzada com a base per implementar aquesta eina.
En el present treball, l’enfocament principal és la classificació de la categoria lèxica de les arrels del mapudüngun reconegudes com a verbals en la font utilitzada per al desenvolupament del sistema d’anàlisi morfològica.
Els resultats d’aquesta revisió categorial beneficien directament l’analitzador computacional, ja que s’implementen tan bon punt es verifiquen. A més, s’espera que aquests resultats ajudin a aclarir algunes incerteses sobre les categories lèxiques en la llengua mapuche.
Aquest treball aborda una tasca preliminar a la identificació de la valència de les arrels verbals, els resultats de la qual es presentaran en un treball posterior que complementa aquest article.
Keywords: Mapuche; Mapudüngun; FST; finite state transducer; morphology; morphosyntax; affixes; suffixes; morphemes; verbs; valency; verbal roots; parts of speech;
Palabras clave: Mapuche; Mapudüngun; FST; transductor de estados finitos; morfología; morfosintaxis; afijos; sufijos; morfemas; verbos; valencia; raíces verbales; categorías gramaticales; categorías léxicas;
Paraules clau: Maputxe; Mapudüngun; FST; transductor d’estats finits; morfologia; morfosintaxi; afixos; sufixos; morfemes; verbs; valència; arrels verbals; categories gramaticals; categories lèxiques;
1 Introduction
This article details the methods used for the reclassification of the lexical roots involved in Mapuche verb themes and presents the results obtained from this process. This task is essential to accurately determine the valency (in a subsequent work, Chandía, 2025) of roots that are genuinely verbal. Without this step, there is a risk of incorrectly attributing verbal properties to roots that do not belong to this lexical category.
The task is described as a reclassification because it involves revisiting the lexical category classification of roots found in Mapudüngun verb themes (see §2.1, p. 2.1) which was originally established by the anthropologist Ineke Smeets [^1] This author will be cited continuously for her 2008 work; thereafter, only her last name and the referenced page will be provided, omitting the year.
in her work on the grammatical description of the Mapuche language, published in 2008 with the title ‘‘A Grammar of Mapuche’’.
This research, like others conducted within the Düngupeyüm [^2] Düngupeyüm is the name of the computational morphological analysis system that has been implemented; the name derives from düngu- ‘language, speech, concern’, -pe- suffix indicating physical or temporal proximity, -yü- suffix indicating a habitual aspect feature, and the affix -m, a nominaliser indicating place or instrument. The translation would be something like ‘instrument always used for language’, i.e., ‘language tool’. The system is publicly available on https://www.chandia.net/dungupeyum
project, relies on the grammatical description provided by Smeets. Our goal is to enhance the computational morphological analysis system that we have developed using finite-state transducers (Chandía, 2012, 2021, 2022), which initially implemented the anthropologist’s work. This focus on the elements and dynamics of the language allows us to refine and expand the capabilities of the system.
For a computational system such as the one described earlier, accurately determining both the lexical category of roots and the valency of verbal roots is crucial for proper analysis. This is because, in Mapudüngun, the addition of suffixes to verbal forms, particularly aspectual ones, depends on the verbal valency or its modification during word formation. Consequently, verbal valency plays a key role in analysing the form to identify the implicated suffixes. Therefore, it was necessary to verify the lexical category of roots involved in the formation of verbal themes, as Mapudüngun naturally verbalises roots from other lexical categories, including nouns, adjectives, adverbs, and even some pronouns and numerals. The aim is to accurately recognise the valency of roots that are genuinely verbal, avoiding misclassification of roots belonging to other categories. This document focuses on the task of lexical classification, while the classification of valency will be addressed in a subsequent study (Chandía, 2025).
During the extensive period dedicated to developing the morphological analysis tool, several observations made by Smeets needed to be verified, and some required modification. She provided a grammatical description of Mapudüngun based on a group of four informants [^3] As asserted by Smeets in email correspondence, the number may be increased by four or five, which where sporadic informants, thus bringing the total to approximately ten.
. As a result, her work does not fully capture the entirety of the Mapuche language or all the phenomena present within it.
The main contribution of Smeets’ ‘‘A Grammar of Mapuche’’, which is particularly valuable for the development of a computational morphological analysis system, is the identification of a prevalent or fixed position for approximately one hundred suffixes that can appear in Mapuche verbal forms. Additionally, she provides the morphotactics [^4] Morphotactics refers to the rules governing the allowed and prohibited combinations of morphemes (both among themselves and with roots) in a language. These rules influence certain forms within the language. For example, in a Mapuche nominal form, morphemes indicating verbal mode cannot appear unless the nominal form has been verbalised. Similarly, in a verbal form, a nominaliser affix and a grammatical person morpheme cannot co-occur. Morphotactic rules can be influenced by grammatical, semantic and morphological factors.
, or combinatorial rules, for most of these suffixes.
In Mapudüngun, a minimal finite verbal form consists of a root (which can be verbal, nominal, adjectival, or adverbial), followed by one or more optional derivational suffixes, and at least one inflectional suffix. For a form to be finite, it must include a mode marker in slot four and a morpheme (which may be null) indicating the person in slot three.
Smeets states that suffixes appear in relatively fixed positions along the verbal form. Based on their placement relative to other suffixes and their function, these verbal affixes are organised into thirty-six numbered slots, arranged from right to left. Slot one is at the far right end of the form, while slot thirty-six is next to the root on the left. Certain slots contain suffixes that are mutually exclusive for grammatical or semantic reasons.
Although it is not unusual to find up to seven or eight suffixes following the root (see E.01), verbs typically contain between four and six suffixes in spontaneous speech.
E.01)
Verb with 10 suffixes (Smeets: 443 .76)
nünieñmarputueyiñmu
  nü            -nie      -ñma  -r      -pu     -tu    -e      -y     -ø   -iñ    -mu
-TV [^5] The meaning of the labels and abbreviations can be found in Appendix §4.1.
.tomar +PRPS +IO   +ITR +LOC +RE +INV +IND +1 +PL +3A
‘they continued to take it away from us’ – ‘continuaron quitándonosla’,
lit.: ‘there they went continuously taking it away from us’ [^6] The Spanish translations of the examples from Smeets, which she translated into English, have been added. Literal interpretations of the Mapuche phrases have also been made in most cases, in an attempt to clarify the meaning of the phrase in the original language.
Figure 1: Graphical representation of the slots in Mapuche verbal forms (root and suffixes) Minimal intransitive verb form of 2nd person plural S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . X X X . . . Minimal transitive verb form of 2nd → 1st persons, plural S . . . . . . . . . . . X . . . . . . . . . . . . . . . . . . X ø X . . . Representation of example E.01 S . . . . X . . . . . X . . . . . . . X X X . . . . . . . . . X . X ø X X
In Figure 1, three Mapuche verbal forms are graphically represented. In this representation, S denotes the verbal theme (stem), and each dot represents a slot. An X indicates a suffix realised in one of the slots, while ø also represents a suffix, specifically a null morpheme. Although the null morpheme has no phonemic or written form, it occupies a corresponding position or slot (Chandía, 2021: 3).
In the Mapuche language, a verbal theme [^7] To simplify, the term “verbal theme” refers to any root, whether simple or complex, to which suffixes are added to create a complete verbal predication. A simple “verbal theme” consists of a single root, while a complex “verbal theme” may involve two roots, a root with some suffixes, or a combination of these elements. Although it could be argued that these are more accurately described as lemmas, they are all referred to as “verbal theme” for the sake of simplicity (Chandía, 2021: 3).
can consist of one, two, or occasionally three roots from different lexical categories, which may combine with each other and sometimes with certain suffixes. A verbal theme can be simple, containing just one root from any lexical category (examples E.02.a, E.02.b, E.02.c), or complex, with several possible configurations:
    A root combined with one or more suffixes (example E.02.d), sometimes up to three suffixes.
    A compound of two roots.
    A less common compound of three roots.
    A compound of two roots with suffixes, typically one but sometimes two or, exceptionally, three suffixes.
Suffixes can follow the first or second root in a compound, or even different suffixes follow each root. Additionally, verbal themes can include a reduplicated root, a root with a reduplicated suffix, or a reduplicated root followed by a suffix.
E.02)
  1. a.
      pichi          -ka        -el (Smeets: 76 .16)
    -AJ.little +CONT +OVN
    ‘[I] still was little (child)’
  2. b.
      düngu       -fi       -nge ! (Smeets: 65 .33)
    -NN.matter +3P +IMP2SG
    ‘talk to him/her’
  3. c.
      tripa         -y      -ø (Smeets: 67 .51)
    -IV.to.leave +IND +3
    ‘he went out / he left’
  4. d.
      küpa         -l     -ün (de Augusta, 1903: 295)
    -IV.to.come +CA +IND1SG
    ‘I brought’; lit.:‘I made come’
1.1 Word Types In Mapudüngun: A State-Of-The-Art Review
The classification of words into lexical types, or the use of a reliable existing classification, is essential in any computational approach to language involving analysis. This is particularly true in our case, which focuses on the morphological analysis of an agglutinative language, where roots of different Parts-of-Speech can generate an inflected verb form through the attachment of affixes.
1.1.1 Dictionaries
To date, the most reliable collections of words classified by their lexical type are the dictionaries published throughout the relatively short written history of the Mapudüngun language, particularly the following three [^8] All three dictionaries are accessible on the CORLEXIM site: https://www.chandia.net/corlexim
(Chandía, 2014).
  1. 1.
    de Valdivia’s dictionary, which reflects a vocabulary less influenced by Spanish, due to its early publication in 1606. A drawback for contemporary studies, however, is that a significant number of words in this dictionary are no longer in use or have undergone varying degrees of semantic and, to some extent, phonological change.
  2. 2.
    Febres’ dictionaries, published in 1765 and 1846 respectively, reflect the influence of Spanish on the Mapudüngun language, resulting from over 300 years of contact between the Mapuche and Spanish cultures. These works often reveal fascinating shifts in meaning; for instance, some terms show notable semantic changes between the first and second dictionaries, or in comparison to Valdivia’s earlier work. In some cases, these changes are intriguing, but at times, they seem disconcerting, as the shifts appear to stem from authorial errors. Nonetheless, Febres’ dictionaries remain valuable resources for contrast and confirmation in cases of uncertainty.
  3. 3.
    de Augusta’s dictionaries, published in 1916, are the most recent in time and offer a reflection of modern Mapudüngun. They clearly show the increased Spanish lexical influence after nearly 500 years of coexistence between the two cultures. This extended period of contact resulted in the imposition of Spanish culture over the Mapuche in almost every aspect of daily life, leading to a process of acculturation that negatively impacted both their culture and their language. The Mapudüngun language, in particular, was prohibited in the Chilean educational system. Despite these challenges, Augusta’s dictionaries remain invaluable due to the remarkable linguistic rigour with which they were compiled.
  4. 4.
    Havestadt’s dictionary is a fourth lexical source occasionally consulted in cases of doubt. Published in 1777, it provides definitions in Latin [^9] Resource available for consultation via the following link: https://benmolineaux.github.io/bookshelf/1777HAVE-Vocab.html. An alternative Spanish version is also available at the following URL: https://benmolineaux.github.io/bookshelf/1777Havestadt-ParsQuarta-MAPU-SPA.html (Martinez and Carmona, 2023).
    and can be a valuable resource, particularly when dealing with variations in forms.
Although these dictionaries are useful, the strong religious bias they carry must be acknowledged. They were compiled by figures such as the Spanish Jesuit de Valdivia, the Catalan Jesuit Febrés, the German Capuchin de Augusta (August Stephan Kathan), and the German Jesuit Havestadt. These works were primarily intended as tools for translating the Gospels and disseminating the Catholic faith, as well as for training priests who would carry out these tasks in Mapuche lands. A second bias stems from the Eurocentric view of the language, typical of earlier studies. This perspective prevailed until relatively recently when typological linguistic studies began to include new points of view that account for the unique features of languages which may differ significantly from the classical grammar models of the Western world.
As mentioned earlier, these dictionaries serve as a reliable source of lexically categorised words. In this regard, the Eurocentric perspective poses only a minor issue for our project, as we employ a broad classification of roots that generally aligns with Universal Grammar models. The system does not utilise the specificity of idiosyncratic subtypes which seems unnecessary for the morphological analysis at this stage; however, they could prove useful for a more fine-grained grammatical analysis in the future.
However, not everything is straightforward. A significant limitation of the Mapuche dictionaries mentioned thus far lies in the way many of the entries are compiled. Often, they are collected with attached suffixes, meaning that the entry –and its lexical categorisation, when provided– corresponds to the derived (and sometimes inflected) form rather than the isolated root. This practice complicates the task of determining the original category of the root, as its meaning –and occasionally its part of speech (PoS)– can shift depending on the suffixes attached. This issue is further addressed in §2.1.
1.1.2 Computational Approach
It is expected that any linguistic computational project considers the classification of the linguistic assets involved. This ensures a robust output from a linguistic perspective. While the system’s functionality may not strictly require such classifications, when dealing with language, the goal is not only for the processes to function correctly but also for the results to be meaningful and useful at a linguistic level.
With this idea in mind, we explored the materials left by one of the few computational projects involving the Mapuche language, conducted in the early 2000s by a prestigious group of researchers from the Language Technologies Institute at Carnegie Mellon University and the Institute of Indigenous Studies at Universidad de La Frontera, supported by the Intercultural Bilingual Education Program of the Chilean Ministry of Education.
AVENUE was the name of this initiative, which produced a document entitled ‘‘Informe Final’’ (Álvarez et al., 2005) summarising the project’s achievements and the works still to be done. When it started in 1999, it was known as NICE (Native-Language Interpretation and Communication Environment), later Avenue. Its goals were to create a simultaneous translation system and to contribute to the preservation and development of Indo-American languages [^10] Avenue worked with three indigenous languages, setting up teams in three countries: Aymara in Bolivia, Quechua in Peru and Mapudüngun in Chile.
. They were committed to creating their own alphabet for the language, transcribing texts according to this alphabet, extracting grammatical rules from the texts, compiling a parallel corpus of Mapudüngun-Spanish, also a glossary, developing a spell-checker of Mapudüngun for OpenOffice, and building a prototype of an automatic translator of Mapudüngun-Spanish. The report mentioned above states that these tasks were carried out at different levels of completion. The section dealing with the automatic translator mentions that it is an example-based machine translation system [^11] Example-based machine translation, or EBMT, works by searching its database for full translations or, more commonly, parts of sentences that have previously been translated by human translators.
. At the time of writing, however, researchers say they are working on a transfer-based machine translation system [^12] In a Transfer-Based Machine Translation (TBMT) system, the original text is first analysed morphologically and syntactically, resulting in a shallow syntactic representation. This representation is then transformed into a more abstract one that emphasises aspects relevant to the translation process and ignores other information. The transfer process converts this latter representation, still linked to the source language, into a representation at the same level of abstraction but linked to the target language. These two representations are called normalised or intermediate representations. From here the process is reversed: the syntactic components generate a representation of the text and finally the translation is generated.
in which the basic constructions of Mapudüngun have been recorded: simple sentences with intransitive and transitive verbs, noun phrases with determiners and modifiers, verb phrases with different temporal and aspectual specifications, passive verb phrases, and verb phrases with inverse agreement [^13] The inversion or reverse form in Mapudüngun is one of the possible expressions of the transitive verb (the others are direct and passive constructions) in which there is a third person acting upon a first or second person, this is expressed by the suffix -e before the mood suffix, and the suffix -(m)ew at the end, provided that the first or second person is a patient; i.e. the order of the constituents, which in Mapudüngun are suffixes, is not changed, but the detailed suffixes are added and combined to change the direction of the action of the verb (Zúñiga, 2006: 114)
. Also relevant is the section dealing with the morphological analyser, which, according to the information in the report, would be an FST system fed by 105 suffixes and some 700 verbs. The document does not specify whether these ‘‘verbs’’ are only verb roots or also other type of roots, such as nominals or adjectivals, nor how the parser works in detail. According to some information that has been gathered, the parser would be developed using the Perl programming language and would also be able to generate sentences in Mapudüngun.
Among the extensive materials, there are indeed lists of lexemes, primarily derived from the dictionaries mentioned in the previous section §1.1.2. Many of these lists adopt the classifications provided in those sources, while others reflect classifications informed by the linguistic knowledge of Mapuche participants and informants involved in the project. However, the project does not clarify the criteria used to establish the parts of speech (PoS) of the listed roots. Consequently, when comparing these lists with the dictionaries, it is impossible to determine which classification is more accurate.
1.1.3 A Lexical Study
In her thesis work (2016), Villena states that her research focuses on examining the authenticity and efficacy of the processes employed in Mapudüngun noun formation. Her objective is to develop criteria for the creation of new nouns in this language. In order to accomplish this objective, she has conducted an analysis of the same dictionaries referenced in §1.1.2.
From the initial statement, it could be inferred that an accurate classification of roots would be of use and performed as a fundamental task. However, the abstract fails to mention that the noun formation under investigation is mainly by derivation processes (compounding is also taken into account), that is to say, by the mechanisms this language commonly uses for the purpose of identifying entities [^14] From our point of view, these are not strictly neologisms, but available (morpho-syntactic) strategies that this language has to deal with the naming of things, new or known. For example, it would not be considered an English neologism: ‘Internet surfer’, even though ‘Internet’ is already a neologism, but ‘Internaut’ would be; it is not registered but possible.
. In any case, the subject of neologicity filters is addressed further down in the work, with the assertion that one of the criteria for determining whether a form is neological is the internal structure of the form (ibid: 11). This suggests that the root component of the neologism should be accurately lexically classified.
However, the prospect of a ‘‘roots classification process’’ appears to have been definitively lost after this note: ‘‘In order to confirm the appropriateness of the lexical categories assigned by Augusta, the most adequate procedure would have been to verify their functioning in discourse. However, the size of the corpus to be analysed made this impossible’’ (ibid: note 82, p. 114). –In the following point §2.1 we explain the procedure followed in our work, which actually implies the process mentioned in Villena’s note–.
Moreover, it appears that accurate lexical classification is not of paramount importance in this research study, particularly in the context of composition, as evidenced by the following citation: ‘‘This finding led to the conclusion that the categories of roots involved in compounding are not crucial in determining the structure of Mapudüngun compounds. Instead, the type of semantic condition maintained between them is deemed to be more significant’’ [^15] As Villena (2016) notes, this is a conclusion of Baker and Fasola (2009) that is supported by the observation of a similarity in the order of the V-N and N-N compounds.
(ibid: 63).
Despite the aforementioned observations, further explorations were conducted on this work, with the objective of identifying any roots that might have undergone a verification of their lexical type. However, ultimately, we were only able to find confirmation of a task that had not been realised, as evidenced by the following few examples:
E.03) (Villena, 2016)
  1. a.
      küdaw-fe             (küdaw- ‘to work’)            ‘worker’
  2. b.
      pu-pülli-amuy      (‘gone into the ground’)       ‘a certain healing art of machi [^16] A machi is the Mapuche chaman.
  3. c.
      yewen                (‘shame’)                         ‘male or female genitalia’
Example E.03.a ((6) b., p.: 53 in the source) shows küdaw- as a verbal root being nominalised by the suffix -fe, which the author says that combines with verbal bases to generate nouns that denote a characteristic agent. But in the dictionaries consulted on the CORLEXIM website (Chandía, 2014) the following entries are found:
  1. küdaw s. The work. (Augusta, F. de (2016a) map-cas)
  2. küdaw N. work. (Smeets, Ineke (2018) map-ing)
  3. küdaw work. (Febrés, A. (1846a) map-cas)
  4. absorto adj. [To be] ngoymaduamkülen (intr.); ngoymakonkülen kiñe küdaw mew. (Augusta, F. de (2016b) cas-map)
As demonstrated above, three authors identify the form as a noun; de Augusta provides an example of the use of the word ‘absorto’ in Mapudüngun (being engrossed in a job), where küdaw is used as a noun without any suffixes, which is in contrast to the classification used by Villena. This example is one of many such instances that can be found in her study.
E.03.b ((92) g., p.: 170 in the source) has been identified as an example of sintagmation, and it is said to represent the Prep-N-V structure. However, a review of dictionaries reveals that:
  1. pu- Vi. to arrive; also puw-. The distribution of pu(w)- is not entirely clear. puw- is most frequent. I found pu-fu-n ‘I arrived, but’ (RR) and iñche pu-n ‘I arrived’ (JM), and I found puw-küle- (RR) only once; müna ayuwüy iñchiu yu puwel ‘he was very glad that we had arrived’; küpa pulen liwen ‘I want to arrive early’; petu puwlaymi küdawmew ‘you are too young to work’; (Smeets, Ineke (2018) map-ing)
Literally translated, E.03.b would mean something like: ‘arrived at the ground as he was going’ or ‘reached the ground on his way’. Perhaps there has been some confusion with another homophonic particle which is a prepositive and which de Augusta and Febrés, among others, define as follows:
  1. pu- artic. pref. pluralises the names of people, animals and some things. pu wentru ‘the men’. pu wangül‘en ‘the stars’. See Augusta (1903, p. 15, 1). (Augusta, F. de (2016a) map-cas)
  2. pu- (part. bef. nouns) denotes plurality; item, at or in; by; under; e.g. pu ovida, sheep (pl.); pu ruka, at home, in the house; pu liwen, in the morning; pu minue, under the ground. (Febrés, A. (1846a) map-cas)
In addition, there is a high probability that the translocative suffix -pu- derives from the same verb, pu-. The affix indicates that the denoted event takes place at a location away from the speaker. This function has also been described by Febrés –and others–:
  1. pu, rpu (interp. to verbs) is to carry out the action of the verb by going from here to there. , v. gr. pepuavin or perpuavin, I’ll drop by to see him; pipuavin or pirpuavin, I’ll pass on to say to him. (Febrés, A. (1846a) map-cas)
E.03.c ((98) b., p.: 176 in the source) is employed to illustrate a case of metonymy, yet it transpires to be a case of semantic calque from the Spanish ’las vergüenzas’. In point of fact, the online DLE dictionary provides the eighth definition for the entry "vergüenza" as follows:
  1. f. pl. Órganos sexuales externos del ser humano. La pintura muestra a Adán con una hoja de parra cubriendo sus vergüenzas. (REAL ACADEMIA ESPAÑOLA: Diccionario de la lengua española, 23.ª ed., [versión 23.8 en línea]. <https://dle.rae.es> [2024-12-21].)
In conclusion, it is evident that the extant literature provides lists of the lexical classification of Mapudüngun roots that are neither complete nor totally reliable. While lexicons represent the most valuable resource in this respect, the remaining works also offer valuable insights that can facilitate the resolution of ambiguities in numerous cases.
2 Verbal roots
As mentioned in the Introduction1, p. 1), for a computational system like the one developed here, correctly identifying the valency of verbal roots is essential for computational morphological analysis. Many of the rules governing the combination of roots and suffixes depend on this classification or on changes in value during the word formation process. This is particularly important for achieving an unambiguous analysis or minimising ambiguity [^17] In a computational morphological analysis system, ambiguity is related to the number of analyses generated by the machine. In other words, the more possible analyses the analyser returns for a lexical form, the more ambiguous the form is. Typically, the sentence context resolves the ambiguity, but this task must be handled by another system: a grammatical disambiguator. However, this tool is only helpful when analysing sentences; for isolated words, or words without context, it is less useful.
. As stated in the introduction, before undertaking the task of valency classification for Mapudüngun verbal roots, a (re)categorisation of the roots that form verbal themes in this language will be carried out. The aim is to more accurately determine the lexical category of these roots, which will help in identifying the roots to be included in the subsequent work on recognising verbal valency (Chandía, 2025). From a computational perspective, this will also enable a better morphological analysis of Mapuche verbs.
2.1 Lexical Category of Roots
The verbal roots listed by Smeets in her dictionary were examined to verify their lexical classification and to modify it if they actually belong to other lexical categories and have been verbalised instead. To achieve this, we primarily referred to the dictionaries by de Augusta, Febrés, and de Valdivia through the CORLEXIM (Chandía, 2014) platform, as well as to the autobiography of the longko [^18] Literally, longko means ‘head’, but similar to other languages, such as Catalan, it is used to denote a chief or leader. In the Mapuche context, a longko is a leader or chief in a broad sense, essentially a community guide who ensures the well-being of his lof (community). On the other hand, a toki is the chief in matters of war. In a more everyday sense, toki means ‘axe’.
Paskwal Koña [^19] The longko Paskwal Koña narrated his life to the Capuchin missionary Ernst Wilhelm von Mösbach, who transcribed it in Mapudüngun. This took place in the 1920s when Koña was still an energetic longko of 80 years. This autobiography is a unique document, not only for the information it provides about the Mapuche language and customs, but also because it is one of the few written sources in Mapudüngun containing words dictated by a Mapuche speaker.
(Wilhelm et al., 1930). Examples of the roots to be verified were sought in all these sources. If any root was found used in isolation —i.e., not as part of a verbal form, nor with affixes attached— it was inferred that the original form is nominal, adjectival, or adverbial, depending on the context and usage. Many of these roots also appear in Smeets’ dictionary with the category listed in the column ‘Final Cat.’ in Table LABEL:tab1. These are non-verbal lexical categories, indicating that these words have been verbalised. In other words, the author records the meaning of the form in its original category as well as in its verbalised version. This column also reflects the lexical category reached after verifying each root in this study. The categories of the roots extracted from Smeets’ dictionary are listed in the ‘Initial Cat.’ column. Some of these roots are recorded by Smeets with an incorporated suffix, which can alter their meaning and valency when the suffix is absent. These suffixes are extracted from the original form and listed in the ‘Extracted suffixes’ column –the following Table LABEL:tab1 is only a sample available for easy verification of what is stated here. The complete table, which contains two hundred and thirty-eight roots, can be found in the Appendix, §4, Table LABEL:tab3, p. LABEL:tab3–.
Table 1: Verbal Roots Reclassified As Other PoS (Sample)
Root Extracted
Suffixes
Initial
Cat.
Final
Cat.
Definition
ad- VI N the visible or characteristic aspects of something or someone, or of a group
chadi- VI N salt
dakel- VT N commission, agreement
[…]
fa- VI Dem this, here
ina- VI Av close
kachu- VI N grass, herb
la- VI Aj dead, deceased
llag- VT Aj half
machi- VI N shaman
nag- VI Av down, below, under, beneath
ngeñi- -ka- VI Aj urgent, pressing
ñochi- VI Aj peaceful, calm
paf- VI N boil, abscess
raküm- VT N closure, wall, protection
tonon- VT N weft
traf- VI Aj together, gathered, joined, fitted, interlocked
[…]
üna- VI N itch
wachi- VT N trap
yafü- VI Aj hard, firm
2.2 Valency Of Verbal Roots
The second step in the classification of roots that participate in Mapuche verbal themes is to specify the valency of verbal roots; as previously mentioned, this task will be presented in Chandía (2025). However, we will preview the results of this study and present a table of some verbal roots that co-occur with the causative suffix -(ü/e)l- (see the list of roots in p. 2.3). We also provide a list of labile roots incorporated into Düngupeyüm and a table of roots that undergo causativisation with the causative suffix -(ü)m- (Table LABEL:tab4).
2.3 Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l-
In the autobiography of the longko (Wilhelm et al., 1930), twenty-one verbal roots [^20] Approximately one hundred and ten verbal roots that were not forming compounds were checked. Since this task took too long, it was decided to postpone the verification of roots when they participate in verbal compounds and exhibit a suffix in one of the forms that the causative can take: -l-, -ül-, -el-.
were found that combine with the causative -l-, as shown in the list bellow. Almost all the roots are intransitive, except for two that form both transitive and intransitive verbs, meaning they are labile or bivalent roots (for a detailed explanation see §3.4 and §3.4.1 in Chandía, 2025). In the following list, the two senses of the labile verbs are marked according to their valency, with -IV (intransitive verb) and -TV (transitive verb).
  1. 1.
    aku- ‘to arrive’
  2. 2.
    allfü- ‘to get injured, hurt, damaged’
  3. 3.
    amu- ‘to go, to go on’
  4. 4.
    apo- ‘to get filled’
  5. 5.
    kümtrü- ‘to rumble, to resonate’
  6. 6.
    küpa- ‘to come’
  7. 7.
    miaw- ‘to prowl, to roam’
  8. 8.
    monge- ‘-IV to live, to heal, to get healed’ / ‘-TV to revive, to heal’
  9. 9.
    montu- ‘to escape’
  10. 10.
[…]
    ‘to wake up’
  11. 11.
    pültrü- ‘to hang’
  12. 12.
    rapi- ‘to vomit’
  13. 13.
    reyü- ‘to mix, to blend’
  14. 14.
    ru- ‘to go along, to go through, to pass’
  15. 15.
    rümu- ‘to sink, to plunge, to immerse oneself’
  16. 16.
    rünga- ‘to embed, to become encrusted’
  17. 17.
    tripa- ‘to leave, to go out, to depart, to set off’
  18. 18.
    ürfi- ‘to drown, to sink’
  19. 19.
    wüda- ‘to divide, to split up, to separate’
  20. 20.
    wüño- ‘to return, to recover’
  21. 21.
    yewe- ‘-IV to feel embarrassed, to be ashamed’ / ‘-TV to respect’
2.4 Collecting Labile Roots In Düngupeyüm
In the Düngupeyüm morphological analysis system, labile roots will now appear in both the transitive and intransitive lists, which inevitably leads to increase ambiguity when analysing verb forms; however, excluding these roots would omit potential analyses. Additionally, the system has been updated to ensure that causatives, whether -(ü)m- or -(ü)l- (Chandía, 2008, §3.3 and §3.4 respectively), do not co-occur with transitive roots. Thus, causatives will only combine with labile roots in their intransitive sense. The following list comprises the labile verb roots that have been incorporated into the system:
List of labile roots collected in the system
  • aye- ‘to laugh’ (-IV)
  • aye- ‘to laugh at; to laugh with’ (-TV)
  • kewa- ‘to fight, to quarrel, to struggle, to combat’ (-IV)
  • kewa- ‘to hit, to strike, to defeat, to overcome’ (-TV)
  • llüka- ‘to get frightened, to be scared, to fear, to be afraid’ (-IV)
  • llüka- ‘to scare, to frighten’ (-TV)
  • meke- ‘to take care of, to deal with, to be busy, to last, to endure’ (-IV)
  • meke- ‘to make’ (-TV)
  • monge- ‘to get life, to heal, to recover’ (-IV)
  • monge- ‘to revive, to bring back to life, to heal’ (-TV)
  • nge- ‘to be’ (-IV)
  • nge- ‘to have’ (-TV) (see nge- -IV ‘to be’, -TV ‘to have’ in Chandía, 2025)
  • püna- ‘to stick, to glue, to adhere’ (-IV)
  • püna- ‘to glue, to paste’ (-TV)
  • waychüf- ‘to fall, to get turned over, to roll, to spin’ (-IV)
  • waychüf- ‘to discard, to flip, to roll, to spin’ (-TV)
  • yewe- ‘to feel embarrassed, to be ashamed’ (-IV)
  • yewe- ‘to respect’ (-TV)
2.5 Verbal Roots That Co-occur With The Causative Suffix -(ü)m-
In Chandía, 2025: §3.3, it will be shown that verifying the presence of the causative suffix -(ü)m- in Mapuche verb forms is a reliable method for determining that the root carrying it is intransitive. However, it is also necessary to verify whether the root is genuinely verbal or it belongs to another lexical category, a task that this document addresses. Furthermore, as noted by Longkon (2011a), Golluscio (2007) and Smeets, not all intransitive roots accept causativisation with -(ü)m-. Therefore, it is important to examine these authors’ claims regarding the role of telicity or atelicity in the selection of the causative suffix. In Table LABEL:tab4, all the roots that have been verified to combine with the causative suffix -(ü)m- are presented, the following Table LABEL:tab2 is only a sample available for easy verification of what is stated in Chandía, 2008: §3.3.
Table 2: Roots That Causativise With -(ü)m- (Sample)
Root Cat. Intransitive sense Root+CA Transitive sense Collected froma
ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
fa Dem this, here fa-m- to be like this K. S.
firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
kon VI to enter, to begin kon-üm- to make come in, to add K. G.
la Aj dead, deceased lang-üm- to kill K. S. G.
(l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
[…]
nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
üy N name üy-üm- to name, to denominate S.
wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
3 Conclusions
In the present work, we undertook a reclassification of the lexical categories of roots involved in verbal themes of the Mapuche language, which are part of the inventory of examples in the descriptive grammar of Mapudüngun developed by the anthropologist Ineke Smeets. This is a preliminary task to valency classification (Chandía, 2025), the study that motivated us to begin this lexical category review. The task of valency classification of Mapuche verbal roots required an accurate classification of the lecixal categories of the roots to avoid attributing verbal features, such as valency, to elements that cannot exhibit them because they are not verbal. A simple verification method for the lexical category was devised, as explained in §2.1, p.2.1, and despite its simplicity, it has proven to be effective, with results shown in Table LABEL:tab3, p. LABEL:tab3).
Next, as the two studies referred to in this article are now complete, we present some of the overall material here, as we have included some of the data generated on completion of both studies in this initial document.
In a computational system like an automatic morphological analyser, where it is crucial to have specific and non-redundant data, ambiguity can be significantly reduced if roots are listed, as much as possible, under a single classification per form. Despite this, we chose to list roots with ambivalent behaviour (see §2.4, p. 2.4) in both lists, intransitive verbal roots and transitive verbal roots, with each list reflecting the root’s valency-related sense. This approach might introduce varying degrees of ambiguity in the analyses produced by the machine. However, besides the ambiguity, it also allows for potential analyses that would be excluded if ambivalent roots were not collected in this manner.
It is essential to first identify the category most likely to be the source of the multiple interpretations or derivations that a form, used as the main component of the verbal theme, might generate (see §2.1, p. 2.1). Therefore, the primary task is to ensure that valency values are sought in inherently verbal roots rather than in verbalised roots derived from other categories (see Table LABEL:tab3, p. LABEL:tab3). This can be challenging in a language like Mapudüngun, as there is often insufficient material or stored knowledge to determine the etymology of much of its vocabulary, largely due to historical and social factors. Techniques like the one employed in this study, while reflecting a highly conditioned perspective, often diverge from fully verifiable data. Nonetheless, it is at least known that Wilhelm et al. (1930) is one of the best references for the Mapuche language (see footnote 19, p. 19), which was the main reason for its use in this research.
To conclude, the decision to publish the study in two parts —lexical classification presented here and valency classification presented in Chandía, 2025— aims to clarify the findings. As stated in the summary, ‘‘The outcomes of this categorical and valency review directly benefit the computational analyser, as they are implemented as soon as they are verified. Additionally, these findings are expected to enhance the understanding of the Mapuche language’’. Readers are invited to review the accompanying article to this first work, which provides a comprehensive understanding of the results of this research.
Acknowledgements
First and foremost, I express my gratitude for the always timely and appropriate support and advice from my thesis supervisors, Irene Castellón and Elizabeth Comelles. Without their guidance, the form and structure of both articles would be far from achieving a good presentation and understandable order, and the terminology used would not always have been accurate.
Secondly, my deepest thanks go to Fernando Zúñiga, who reviewed and corrected not only the content aspects of both articles but also some formal aspects. Without his extensive linguistic knowledge in general and of Mapudüngun in particular, many of the points discussed along the two articles would have been nothing more than imprecise approximations to the linguistic ideas and situations of the Mapuche language addressed in both works.
4 Appendix
4.1 Labels And Abbreviations
+1   First person
+1t2A   First agent to Second patient
+2   Second person
+3   Third person
+3A   Third person agent
+3P   3rd person patient (Differential Object Marker)
+ADJDO   Doable action adjectiviser
Aj, -AJ   Adjective
-AP   Anaphora
Av, -AV   Adverb
+BEN   Benefactive
CA, +CA   Causative
Conj, -CJ   Conjunction
-CONT   Continuative
-CR.IV   Intransitive compound verbal stem
-CR.TV   Transitive compound verbal stem
Dem, -DP   Demonstrative
+DL   Dual number
+EXP   Experimentative
+FAC   Factual
+FORCE   Obligative
G.   Lucía Golluscio
+FUT   Future
+HAB   Habituative
+IND   Indicative
+IND1SG   Indicative 1st person singular
+INST   Instrumentaliser
Int, -IP   Interrogative
+INV   Transitive inversion marker
+IO   Indirect object
+ITR   Interruptive
+IVN   Instrumental verbal noun
K.   Paskwal Koña
+LOC   Locative
+MIO   More implied object
N, -NN   Noun (substantive)
+NEG   Negator
+NOM   Nominaliser
+NU   Numeral
+$O$   Oblique object
+OVN   Objective verbal noun
+PASS   Passive voice
+PFPS   Perfect persistent
+PL   Plural number
+PLR   Pluraliser
+PRPS   Progressive persistent
+PVN   Plain verbal noun
+RE   Repetitive/Restaurative
+REF   Reflexive/Reciprocal
+RI   Ruptured implicature
-RNNR   Reduplicated nominal root
S.   Ineke Smeets
+SFR   Stem formant
+SG   Singular number
+SJI   Subjunctive affix in imperatives forms
-SP   Possessive pronoun
+ST   Estative
+SVN   Subjective verbal noun
+TH   Thither
+TR   Transitiviser
VI, -IV   Intransitive verb
+VRB   Verbaliser
VT, -TV   Transitive verb
4.2 Tables
Table 3: Verbal Roots Reclassified As Other PoS
Root Extracted
Suffixes
Initial
Cat.
Final
Cat.
Definition
ad- VI N the visible or characteristic aspects of something or someone, or of a group
af- VI N end
afü- VI Aj cooked, ripe
allush- VI Aj lukewarm
alü- VI Av very, much
angim- VT N jerky
angka- VI N middle, half
antü- VI N Sun, day, time
are- VI Aj hot, warm
ariñ- VI Aj burnt, smoked
arof- VI N sweat
asul- VI Aj blue
awi- VI Aj hot
awka- VI Aj rebellious, wild
aylen- VI N ember(s)
aywiñ- VI N shade
chadi- VI N salt
chafo- VI N cough
chali- VI N greeting
challwa- VI N fish
chape- VT N braid (in the hair)
che- VI N person, people
chedkuy- VT N father/son in law (of a man)
chilla- VT N saddle, chair
chod- VI Aj yellow
chum- VT Int how, what
chüngküd- VI Aj circular, rounded
dakel- VT N commission, agreement
dew- VI Av already, since, because, no more, only, finally
domo- VI N woman
[…]
duam- VT N matter, issue, need, intention
dumdum- VI N penumbra
düngu- VT N issue, matter, subject, concern, word, language
[…]
elfal- VT N commission
eñum- VI Aj comfortingly/pleasantly warm
epew- VI N tale, story, history, narration
fa- VI Dem this, here
fe- VI Dem that, there
fente- VT Av so, much
fentre- VI Av very, much
fey- VI Dem this, that, she, he
filla- VI N scarcity, lack
fitrun- VI N smoke
fücha- VI Aj big, old
fülang- VI Aj white
funa- VI Aj rotten
füre- VI Aj spicy, hot
futrül- VT N heap, pile
füw- VT N thread, wool
ina- VI Av close
ingka- VT N defender
kachu- VI N grass, herb
kafe- VI N coffee
kallfü- VI Aj blue, purple, violet
kangka- VT Aj roasted
karü- VI Aj green
kashü- VI Aj gray
katrü- VT Aj cut off, interrupted
kawe- VI N oar
kellu- VT N help
kelü- VI Aj red
kim- VT Aj wise, knower
kochü- VI Aj sweet
kofke- VT N bread
kolü- VI Aj rust, reddish-brown (colour)
kompañ- VT N partner
koñoll- VI Aj purple, violet
koñü- VI N baby, newborn
korü- VI N soup, broth
kotrü- VI Aj salty, sour, bitter
kowkow- VI N owl
küdaw- VI N work
kude- VI N bet
[…]
küllay- -tu- VT N quillaya, quillay soap
külü- VI Aj tilted, slanted
küme- VI Aj good
kunaw- VI Aj inflated, floating
kupül- VT Aj fixed, secure
kuram-b VI N egg
küri- VI Aj lustful, lecherous
kurü- VI Aj black
kürüf- VI N wind
kushe- VI Aj elderly, old (woman)
kutran- VI N illness, disease, pain, suffering
kütrüng- VT N bundle, bunch
küyen- VI N moon, month
la- VI Aj dead, deceased
lantra- VI Aj bulging, thick
lawen- VT N medicine (medicinal herb)
lef- VI Aj fast (going/moving)
leg- VI Aj certain, accurate, correct
lichi- VI N milk
lif- VI Aj clean
[…]
llag- VT Aj half
llekü- VI Av close
llitu- VI Aj initial, first
[…]
llüngüd- VT Aj flat, level
llüpañ- VT Aj lying, reclining
lüf- VI Aj scorched, roasted
lüp-d -üm- VT Aj scorched, roasted
lüykü- VI N drop
machi- VI N shaman
[…]
mafü- VT N dowry
[…]
mangkü- VT N kick, boot
mañum- VT N gratitude, thankfulness
mawün- VT N rain
may- VT Conj yes
me- VI N excrement, scum, dross
melkay- VI Aj slippery
[…]
minggako- VI N communal/community work
misha- VT N dining companion
[…]
motri- VI Aj fat
moy- VI N boil
moyo- VT N breast
muday- VI N chicha
[…]
müna- -le- VI Av many, much
mupi- VI Aj true, genuine
müpü- VI N wing
[…]
mütrüm- VT N call
nag- VI Av down, below, under, beneath
[…]
ngeñi- -ka- VI Aj urgent, pressing
ngoyma- VT Aj unproductive, neglected
ngüdi- VT N obstruction
ngülam- VT N advice, tip
ngüñü- VI N hunger
nor- VI Aj straight (in a physical or figurative sense), correct
notu- -ka- VT N obstinacy, blindness
nüla- VT Aj open
nümü- VI N smell, odour
[…]
nürüf- VT Aj closed
nütram- VT N recount, conversation, story
nüyün- VI N tremor, quake
ñochi- VI Aj peaceful, calm
ñum- VI N upper millstone
paf- VI N boil, abscess
[…]
pali- VI N paline ball
panü- VT N fathom
papel- VT N paper
[…]
perimo- -ntu- VI N omen
perkü- VI N fart
pewma- VI N daydream, dream
pichi- $similar-to\sim püchü- simi$$similar-to\sim püt$ü- VI Aj little
pidku- VT Aj stew (referring to a dish typically involving legumes cooked in a pot)
piku- VI N North
pilu- VI Aj deaf
pire- VI N snow, hail
[…]
pitru- VI N scabies
pod- VI Aj dirty
pofre- VI Aj poor
poy- VI N tumour, swelling
püd- VI Aj thick, dense, condensed
pülle- VI Av close, next
püntü- VI N piece, chunk
püñeñ- VI N daughter or son (to the mother or her sister)
püro- VT N knot
pütre- VT Aj blazing, scorching, burning, hot
pütrü- VI Av in abundance, a lot, plentifully
raküm- VT N closure, wall, protection
rangiñ- VI N middle, half
re- -le- VI Aj alone, exclusive, pure
riku- VI Aj rich (it may be related to rükü- (further down)
ruka- VT N house
rükü- VI Aj mean, petty, greedy, miserly, tight-fisted, stingy
rümpel- VI Aj jealous
[…]
rutra- -tu- VT N pinch
tonon- VT N weft
traf- VI Aj together, gathered, joined, fitted, interlocked
traf-iya-f VI Av last night, since last night’s fall
tralka- VI N thunder, shotgun
tran- VI Aj fallen (especially referring to trees)
trangliñ- VI N frost
trapel- VT N cord, string
trar- VI N pus
[…]
traytray- VI N gargle, or similar gargling sound
trem- VI Aj grown, adult, mature
tremo- VT Aj healthy, robust, splendid
trintri- VI Aj curly
tripantu- VI N year
tromü- VI N cloud
tror- VI N foam
trüfon- VI N cough
trufür- VI N dust
[…]
trülke- VT N leather
[…]
trutruka- VI N a wind instrument like a long cornet
tute- VT Aj correct, accurate
üküm- VI N silence
ül- VI N song, singing
[…]
üna- VI N itch
üng- -üm- VT AV while
[…]
üyag- -tu- VT Aj both
wachi- VT N trap
waw- VI N valley, lowland, low-lying area
wechod- VI N hole
wedwed- VI Aj mischievous
welu- VI Aj alternative, alternating
wentru- VI N man
wew- VT N victory, triumph, gain, profit
[…]
wif- VI Aj wide, broad
[…]
wilüf- VI Aj shiny
wim- VI Aj tame, accustomed
wima- VT N rod, staff
witral- VT N warp, loom
wün- VI N mouth, dawn, sunrise
wüne- VI Aj fore, anterior, superior
wütre- VI Aj cold
yafü- VI Aj hard, firm
yall- VT N son
[…]
yochi- VI Aj tight, fitted
yung- -üm- VT Aj pointed, sharp
a. Küllay is a tree endemic to the central region of Chile; scientific name Quillaja saponaria, also known as ‘soapbark tree’.
b. In conjugated (verbalised) forms, the root kuram ‘egg’ is alternatively analysed as -NN.kürastone+CA.m34… by Düngupeyüm, that is to say, as kura-m, a form that may be interpreted as ‘caused to be a stone’, which may well be the origin of kuram ‘egg’.
[…]
d. The causative suffix -(ü)m- triggers a radical change in some roots, causing those ending in f to shift to p (see Chandía, 2025, E.06), indicating that this root (lüp-) is the same as the one above (lüf-).
e. Palin is a sport similar to field hockey, referred to as chueca by Spaniards and Creoles.
[…]
Table 4: Roots That Causativise With -(ü)m-
Root Cat. Intransitive sense Root+CA Transitive sense Collected froma
1. ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
2. afü Aj cooked, ripe afü-m- to cook K. S. G.
3. ali VI to get heated up ali-m- to heat up S.
4. angkü VI to get dry (up), to be sterile, to be arid angkü-m- to dry (up) K. S.
5. anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
6. are Aj hot, warm are-m- to warm up S. G.
7. chaw N father, man chaw-üm-b to incubate K. S.
8. che N person, people, human being che-m- to relatec K.
9. chong VI to go out, to extinguish, to dry up chong-üm- to turn off, to extinguish, to dry S.
10. chu Int interrogative particle chu-m to ask a question d K. S.
11. fa Dem this, here fa-m- to be like this K. S.
12. fay VI to ferment fay-üm- to [make] ferment S.
13. fe Dem that, there fe-m- to be like that K. S.
14. fey VI to get adapted, to fit fey-üm- to adapt, to fit, to insert S.
15. firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
16. fül Av close fül-üm- to bring closer, to approximate, to bring near K. S.
17. füy VI to squeeze (itself), to hold on (to itself) füy-üm- to tighten, to hold K.
18. illku VI to get angry illku-m- to contradict, to make angry K.
19. kofi VI to get heat up, to get overheat kofi-m- to heat, to overheat S.
20. kon VI to enter, to begin kon-üm- to make come in, to add K. G.
21. kudu VI to lie (down), to go to bed kudu-m- to lay out K.
22. küme Aj good, nice, pleasant küme-m- to add value, to improve K.
23. kümpu N piece, chunk, bit kümpu-m- to chop, to cut into pieces, to tear apart, to shred K.
24. küpa VI to come küpa-m- to bring, to introduce K.
25. la Aj dead, deceased lang-üm- to kill K. S. G.
26. laf Aj flat, level lap-üm- to stretch, to extend S.
27. law VI to lose hair/feathers law-üm- to shave, to shear, to pluck S.
28. lef Aj fast (on going or moving) lep-üm- to make run (animals), to make flee, to make turn away S. G.
29. (l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
30. lif Aj clean lip-üm- to clean K.
31. llangkü VI to fall (down) llangkü-m- to drop, to lose K. S.
32. lleg VI to grow (up), to come up llek-üm- to raise, to cultivate, to grow (plants), to bring out S. G.
[…]
34. llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
35. lüf Aj burned, roasted p-üm- to burn, to set fire, to roast S. G.
36. lüg Aj white lüg-üm- to whiten, to bleach, to lighten (in color) S.
37. lulu N noise, crash, upheaval lulu-m- to make noise, to cause crashing K.
38. nag Av down, underneath nak-üm- to go down, to lower, to descend K. S. G.
[…]
40. neng VI to move (oneself) neng-üm- to move, to shake, to shudder, to direct K.
41. neyküe VI to unleash, to get free neykü-m- to release, to let go, to free S.
42. ngül VI to get gathered, to come together, to join ngül-üm- to gather, to bring together, to collect, to harvest K. S.
43. nong VI to get broken down, to get damaged, to get scorched nong-üm- to spoil, to damage, to scorch K.
44. nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
45. ñam VI to get lost, to get astray, to get loose ñam-üm- to lose, to release, to let go, to make disappear S. G.
46. ñif VI to dry out ñip-üm- to dry S.
47. ñom Aj meek, gentle, tame, docile, obedient ñom-üm- to tame, to domesticate, to tame K. S.
48. pelof N light, clarity pelo-m- to illuminate, to light up, to turn on the light S.
49. pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
50. piwü VI to dry out piwü-m- to dry (up) K. S.
51. pod N dirt, filth, uncleanliness pod-üm- to dirty, to soil S.
52. poñü N potato poñü-m- feed with potatoes K.
53. püd Aj thick, dense, scattered, dispersed püd-üm- to spread, to scatter, to thicken, to extend, to spread out, to stretch, to ration, to distribute K. S.
54. pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
55. püna VI to get stuck, to get glued, to get adhered püna-m- to stick, to glue, to adhere S.
56. püra VI to ascend, to go up, to climb püra-m-g to lift, to raise, to pick up, to keep in mind, to remember, to make evident, to highlight K. S. G.
57. püsha VI to run out, to be finished püsha-m- to finish, to end, to spend, to use up, to exhaust G.
58. pütre Aj burning, red-hot pütre-m- to smoke, to fumigateh K.
59. puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
60. rüngü VI to be pulverised, to be ground, to be crushed rüngü-m- to grind, to mill, to crush S.
61. traf Aj together, joined, gathered, fitted trap-üm- to gather, to collect, to put together, to grab, to catch, to pick up K. S.
62. tral VI to get hit, to get bumped, to get dented, to get bruised, to collide, to crash, to clash tral-üm- to hit, to strike, to beat K.
63. tran Aj fallen (especially referring to trees) tran-üm- to make fall, to knock down, to knock over, to topple, to lay down G.
64. trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
65. trem Aj grown, adult, mature trem-üm- to raise, to breed K. S. G.
66. trof VI to burst, to break, to crack, to snap trop-üm- to split, to break S.
67. trong Aj thick, dense trong-üm- to roof (houses)i K.
68. trür Aj even, level, tight, composed trür-üm- to equalise, to compensate, to balance, to indemnify K. S.
69. tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
70. üng Av while üng-üm- to wait K.
[…]
72. ürkü VI to get tired ürkü-m- to tire, to tire out G.
[…]
74. üy N name üy-üm- to name, to denominate S.
75. wadkü VI to boil wadkü-m- to boil, to bring to a boil K. S. G.
76. wallü VI to get surrounded, to get encompassed, to get completed wallü-m- to tell or recount in detail K.
77. wef VI to appear, to show itself wep-üm- to make appear, to find, to show G.
78. welli VI to get empty welli-m- to empty, to vacate K. S.
79. wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
80. wüda VI to divide, to split up wüda-m- to divide, to distribute, to share, to split K. S.
81. wünü VI to get stretched, to get elongated wünü-m- to stretch, to lengthen S.
82. yall N son (of a man) yall-üm- to multiply, to engender S.
83. yewe VI to feel embarrassed, to be ashamed yewe-m- to embarrass G.
84. yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
a. In the ‘‘Collected from’’ column, the reference work containing examples with the corresponding root is identified. K. stands for Paskwal Koña in Wilhelm et al., 1930; S. for Smeets; and G. for Golluscio, 2007.
b. The Mapuche use the term chaw-üm- ‘fathering’, or more literally ‘to cause to father’, to refer to ‘incubating eggs’.
c. In reality, to convey the meaning of ‘to relate’, the compound che-m-ye- needs to be formed, which is analysed as -NN.che_person_people+CA.m34+TV.ye_to_carry. This is the way to express a relationship with someone in Mapudüngun. For example, laku-ye-fi-ñ translates to ‘he is my paternal grandfather’, which literally means ‘I carry him as a grandfather’ (Becerra, 2011: 115).
d. In the case of interrogatives like ‘‘what, which, how’’, etc., staring with chu-, these are formed by adding other suffixes. For example, chu-m-al means ‘what for’, chu-m-nge-chi means ‘how’, chu-m-ül means ‘when’, and so on.
e. It is very likely that neykü is a lexicalised form, composed of the root nel- listed in line 39, and the unproductive suffix -kü-, which is found in a few verbal roots but has not been documented or described in existing literature.
f. pelo is a lexicalised form of a nominalised verb (see example E.28.a in Chandía, 2025).
g. The form püra-m- generates compounds with many action verbs to indicate an upward direction, such as in tralkatunpüramün, which means ‘to shoot into the air’. It is also used figuratively, as in kimpüramün, meaning ‘to become aware’; literally: ‘to raise knowledge’.
h. When pütre is nominalised as pütrem -AJ.pütre_blazing+CA.m34+NOM.ø, it means ‘cigar’ or ‘tobacco’.
i. trongum- is to thatch a ruka (a traditional Mapuche house) with reed or tall grass, these materials must be placed very densely, ensuring a good level of thickness.

### v3

@@
 This work addresses a preliminary task to identify the valency of true verbal roots, the results of which will be presented in a subsequent work that complements this article.
-Resumen
+?abstractname?
 Tras haber desarrollado un sistema computacional para el análisis morfológico de la lengua Mapuche y de evaluarlo con textos de diversos autores y estilos, se ha visto la necesidad de verificar los postulados lingüísticos de la fuente utilizada como base para implementar esta herramienta.
@@
 Este trabajo aborda una tarea preliminar a la identificación de la valencia de las raíces verbales, cuyos resultados se presentarán en un trabajo posterior que complementa este artículo.
-Resum
+?abstractname?
 Després d’haver desenvolupat un sistema computacional d’anàlisi morfològica de la llengua Mapuche, i d’avaluar-lo amb textos de diversos autors i estils, es va veure la necessitat de verificar els postulats lingüístics de la font utilitzada com a base per implementar aquesta eina.
@@
 Aquest treball aborda una tasca preliminar a la identificació de la valència de les arrels verbals, els resultats de la qual es presentaran en un treball posterior que complementa aquest article.
-Keywords: Mapuche; Mapudüngun; FST; finite state transducer; morphology; morphosyntax; affixes; suffixes; morphemes; verbs; valency; verbal roots; parts of speech;
-Palabras clave: Mapuche; Mapudüngun; FST; transductor de estados finitos; morfología; morfosintaxis; afijos; sufijos; morfemas; verbos; valencia; raíces verbales; categorías gramaticales; categorías léxicas;
-Paraules clau: Maputxe; Mapudüngun; FST; transductor d’estats finits; morfologia; morfosintaxi; afixos; sufixos; morfemes; verbs; valència; arrels verbals; categories gramaticals; categories lèxiques;
+Keywords: Mapuche; Mapudüngun; FST; finite state transducer; morphology; morphosyntax; affixes; suffixes; morphemes; verbs; valency; verbal roots; parts of speech
+Palabras clave: Mapuche; Mapudüngun; FST; transductor de estados finitos; morfología; morfosintaxis; afijos; sufijos; morfemas; verbos; valencia; raíces verbales; categorías gramaticales; categorías léxicas
+Paraules clau: Maputxe; Mapudüngun; FST; transductor d’estats finits; morfologia; morfosintaxi; afixos; sufixos; morfemes; verbs; valència; arrels verbals; categories gramaticals; categories lèxiques
 1   Introduction
 This article details the methods used for the reclassification of the lexical roots involved in Mapuche verb themes and presents the results obtained from this process.
-This task is essential to accurately determine the valency (in a subsequent work, Chandía, 2025) of roots that are genuinely verbal.
+This task is essential to accurately determine the valency (in a subsequent work, Chandía, 2026) of roots that are genuinely verbal.
 Without this step, there is a risk of incorrectly attributing verbal properties to roots that do not belong to this lexical category.
-The task is described as a reclassification because it involves revisiting the lexical category classification of roots found in Mapudüngun verb themes (see §2.1, p.
-2.1) which was originally established by the anthropologist Ineke Smeets [^1] This author will be cited continuously for her 2008 work; thereafter, only her last name and the referenced page will be provided, omitting the year.
-in her work on the grammatical description of the Mapuche language, published in 2008 with the title ‘‘A Grammar of Mapuche’’.
-This research, like others conducted within the Düngupeyüm [^2] Düngupeyüm is the name of the computational morphological analysis system that has been implemented; the name derives from düngu- ‘language, speech, concern’, -pe- suffix indicating physical or temporal proximity, -yü- suffix indicating a habitual aspect feature, and the affix -m, a nominaliser indicating place or instrument.
+The task is described as a reclassification because it involves revisiting the lexical category classification of roots found in Mapudüngun verb themes (see §2.1) which was originally established by the anthropologist Ineke Smeets in her work on the grammatical description of the Mapuche language, published in 2008 with the title “A Grammar of Mapuche”.
+This research, like others conducted within the Düngupeyüm [^1] Düngupeyüm is the name of the computational morphological analysis system that has been implemented; the name derives from düngu- ‘language, speech, concern’, -pe suffix indicating physical or temporal proximity, -yü suffix indicating a habitual aspect feature, and the affix -m, a nominaliser indicating place or instrument.
 The translation would be something like ‘instrument always used for language’, i.e., ‘language tool’.
-The system is publicly available on https://www.chandia.net/dungupeyum
-project, relies on the grammatical description provided by Smeets.
+The system is publicly available on https://www.chandia.net/dungupeyum/ project, relies on the grammatical description provided by Smeets (2008).
 Our goal is to enhance the computational morphological analysis system that we have developed using finite-state transducers (Chandía, 2012, 2021, 2022), which initially implemented the anthropologist’s work.
@@
 The aim is to accurately recognise the valency of roots that are genuinely verbal, avoiding misclassification of roots belonging to other categories.
-This document focuses on the task of lexical classification, while the classification of valency will be addressed in a subsequent study (Chandía, 2025).
-During the extensive period dedicated to developing the morphological analysis tool, several observations made by Smeets needed to be verified, and some required modification.
-She provided a grammatical description of Mapudüngun based on a group of four informants [^3] As asserted by Smeets in email correspondence, the number may be increased by four or five, which where sporadic informants, thus bringing the total to approximately ten.
-. As a result, her work does not fully capture the entirety of the Mapuche language or all the phenomena present within it.
-The main contribution of Smeets’ ‘‘A Grammar of Mapuche’’, which is particularly valuable for the development of a computational morphological analysis system, is the identification of a prevalent or fixed position for approximately one hundred suffixes that can appear in Mapuche verbal forms.
+This document focuses on the task of lexical classification, while the classification of valency will be addressed in a subsequent study (Chandía, 2026).
+During the extensive period dedicated to developing the morphological analysis tool, several observations made by Smeets (2008) needed to be verified, and some required modification.
+She provided a grammatical description of Mapudüngun based on a group of four informants [^2] As asserted by Smeets in email correspondence, the number may be increased by four or five, which where sporadic informants, thus bringing the total to approximately ten..
+As a result, her work does not fully capture the entirety of the Mapuche language or all the phenomena present within it.
+The main contribution of Smeets (2008)’ “A Grammar of Mapuche”, which is particularly valuable for the development of a computational morphological analysis system, is the identification of a prevalent or fixed position for approximately one hundred suffixes that can appear in Mapuche verbal forms.
 Additionally, she provides the morphotactics [^3] Morphotactics refers to the rules governing the allowed and prohibited combinations of morphemes (both among themselves and with roots) in a language.
@@
 Similarly, in a verbal form, a nominaliser affix and a grammatical person morpheme cannot co-occur.
-Morphotactic rules can be influenced by grammatical, semantic and morphological factors.
-, or combinatorial rules, for most of these suffixes.
+Morphotactic rules can be influenced by grammatical, semantic and morphological factors., or combinatorial rules, for most of these suffixes.
 In Mapudüngun, a minimal finite verbal form consists of a root (which can be verbal, nominal, adjectival, or adverbial), followed by one or more optional derivational suffixes, and at least one inflectional suffix.
 For a form to be finite, it must include a mode marker in slot four and a morpheme (which may be null) indicating the person in slot three.
-Smeets states that suffixes appear in relatively fixed positions along the verbal form.
+Smeets (2008) states that suffixes appear in relatively fixed positions along the verbal form.
 Based on their placement relative to other suffixes and their function, these verbal affixes are organised into thirty-six numbered slots, arranged from right to left.
@@
 Although it is not unusual to find up to seven or eight suffixes following the root (see E.01), verbs typically contain between four and six suffixes in spontaneous speech.
-E.01)
-Verb with 10 suffixes (Smeets: 443 .76)
-nünieñmarputueyiñmu
-  nü            -nie      -ñma  -r      -pu     -tu    -e      -y     -ø   -iñ    -mu
--TV [^5] The meaning of the labels and abbreviations can be found in Appendix §4.1.
-.tomar +PRPS +IO   +ITR +LOC +RE +INV +IND +1 +PL +3A
-‘they continued to take it away from us’ – ‘continuaron quitándonosla’,
-lit.: ‘there they went continuously taking it away from us’ [^6] The Spanish translations of the examples from Smeets, which she translated into English, have been added.
-Literal interpretations of the Mapuche phrases have also been made in most cases, in an attempt to clarify the meaning of the phrase in the original language.
-Figure 1: Graphical representation of the slots in Mapuche verbal forms (root and suffixes) Minimal intransitive verb form of 2nd person plural S .
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-X X X .
-.
-.
-Minimal transitive verb form of 2nd → 1st persons, plural S .
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-X .
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-.
-X ø X .
-.
-.
-Representation of example E.01 S .
-.
-.
-.
-X .
-.
-.
-.
-.
-X .
-.
-.
-.
-.
-.
-.
-X X X .
-.
-.
-.
-.
-.
-.
-.
-.
-X .
-X ø X X
+E.01 10 suffixes Mapuche verb nünieñmarputueyiñmu
+  1. nü      -nie   -ñma   -r    -pu    -tu   -e    -y    -ø   -iñ   -mu (Smeets, 2008: 443; 76)
+    tv.take +prps +io +itr +loc +re +inv +ind +1 +pl +3a
+    ‘they continued to take it away from us’; Literal: ‘there they went continuously taking it away from us’
+Minimal intransitive verb form of 2nd person plural
+S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . X X X . . .
+Minimal transitive verb form of 2nd → 1st persons, plural
+S . . . . . . . . . . . X . . . . . . . . . . . . . . . . . . X ø X . . .
+Representation of E.01
+S . . . . X . . . . . X . . . . . . . X X X . . . . . . . . . X . X ø X X
+?figurename? 1: Graphical representation of the slots in Mapuche verbal forms (root and suffixes)
 In Figure 1, three Mapuche verbal forms are graphically represented.
@@
 can consist of one, two, or occasionally three roots from different lexical categories, which may combine with each other and sometimes with certain suffixes.
-A verbal theme can be simple, containing just one root from any lexical category (examples E.02.a, E.02.b, E.02.c), or complex, with several possible configurations:
-    A root combined with one or more suffixes (example E.02.d), sometimes up to three suffixes.
+A verbal theme can be simple, containing just one root from any lexical category (E.02a, E.02b, E.02c), or complex, with several possible configurations:
+  • A root combined with one or more suffixes (E.02d), sometimes up to three suffixes.
   • A compound of two roots.
@@
 Additionally, verbal themes can include a reduplicated root, a root with a reduplicated suffix, or a reduplicated root followed by a suffix.
-E.02)
+E.02
   1. a.
-      pichi          -ka        -el (Smeets: 76 .16)
-    -AJ.little +CONT +OVN
+    pichi      -ka     -el (Smeets, 2008: 76; 16)
+    aj.little +cont +ovn
     ‘[I] still was little (child)’
   2. b.
-      düngu       -fi       -nge ! (Smeets: 65 .33)
-    -NN.matter +3P +IMP2SG
+    düngu     -fi   -nge ! (Smeets, 2008: 65; 33)
+    nn.speach +3p +imp2sg
     ‘talk to him/her’
   3. c.
-      tripa         -y      -ø (Smeets: 67 .51)
-    -IV.to.leave +IND +3
+    tripa     -y    -ø (Smeets, 2008: 67; 51)
+    iv.leave +ind +3
     ‘he went out / he left’
@@
     küpa    -l    -ün (de Augusta, 1903: 295)
-    -IV.to.come +CA +IND1SG
-    ‘I brought’; lit.:‘I made come’
+    iv.come +ca +ind1sg
+    ‘I brought’; Literal: ‘I made come’
 1.1   Word Types In Mapudüngun: A State-Of-The-Art Review
@@
 1.1.1   Dictionaries
-To date, the most reliable collections of words classified by their lexical type are the dictionaries published throughout the relatively short written history of the Mapudüngun language, particularly the following three [^8] All three dictionaries are accessible on the CORLEXIM site: https://www.chandia.net/corlexim
-(Chandía, 2014).
+To date, the most reliable collections of words classified by their lexical type are the dictionaries published throughout the relatively short written history of the Mapudüngun language, particularly the following three [^5] All three dictionaries are accessible on the CORLEXIM site: https://www.chandia.net/corlexim (Chandía, 2014).:
   1. 1.
@@
 With this idea in mind, we explored the materials left by one of the few computational projects involving the Mapuche language, conducted in the early 2000s by a prestigious group of researchers from the Language Technologies Institute at Carnegie Mellon University and the Institute of Indigenous Studies at Universidad de La Frontera, supported by the Intercultural Bilingual Education Program of the Chilean Ministry of Education.
-AVENUE was the name of this initiative, which produced a document entitled ‘‘Informe Final’’ (Álvarez et al., 2005) summarising the project’s achievements and the works still to be done.
+AVENUE was the name of this initiative, which produced a document entitled “Informe Final” (Álvarez et al., 2005) summarising the project’s achievements and the works still to be done.
 When it started in 1999, it was known as NICE (Native-Language Interpretation and Communication Environment), later Avenue.
-Its goals were to create a simultaneous translation system and to contribute to the preservation and development of Indo-American languages [^10] Avenue worked with three indigenous languages, setting up teams in three countries: Aymara in Bolivia, Quechua in Peru and Mapudüngun in Chile.
-.
+Its goals were to create a simultaneous translation system and to contribute to the preservation and development of Indo-American languages [^7] Avenue worked with three indigenous languages, setting up teams in three countries: Aymara in Bolivia, Quechua in Peru and Mapudüngun in Chile..
 They were committed to creating their own alphabet for the language, transcribing texts according to this alphabet, extracting grammatical rules from the texts, compiling a parallel corpus of Mapudüngun-Spanish, also a glossary, developing a spell-checker of Mapudüngun for OpenOffice, and building a prototype of an automatic translator of Mapudüngun-Spanish.
 The report mentioned above states that these tasks were carried out at different levels of completion.
-The section dealing with the automatic translator mentions that it is an example-based machine translation system [^11] Example-based machine translation, or EBMT, works by searching its database for full translations or, more commonly, parts of sentences that have previously been translated by human translators.
-.
+The section dealing with the automatic translator mentions that it is an example-based machine translation system [^8] Example-based machine translation, or EBMT, works by searching its database for full translations or, more commonly, parts of sentences that have previously been translated by human translators..
 At the time of writing, however, researchers say they are working on a transfer-based machine translation system [^9] In a Transfer-Based Machine Translation (TBMT) system, the original text is first analysed morphologically and syntactically, resulting in a shallow syntactic representation.
@@
 in which the basic constructions of Mapudüngun have been recorded: simple sentences with intransitive and transitive verbs, noun phrases with determiners and modifiers, verb phrases with different temporal and aspectual specifications, passive verb phrases, and verb phrases with inverse agreement [^10] The inversion or reverse form in Mapudüngun is one of the possible expressions of the transitive verb (the others are direct and passive constructions) in which there is a third person acting upon a first or second person, this is expressed by the suffix -e before the mood suffix, and the suffix -(m)ew at the end, provided that the first or second person is a patient; i.e.
-the order of the constituents, which in Mapudüngun are suffixes, is not changed, but the detailed suffixes are added and combined to change the direction of the action of the verb (Zúñiga, 2006: 114)
-.
+the order of the constituents, which in Mapudüngun are suffixes, is not changed, but the detailed suffixes are added and combined to change the direction of the action of the verb (Zúñiga, 2006: 114)..
 Also relevant is the section dealing with the morphological analyser, which, according to the information in the report, would be an FST system fed by 105 suffixes and some 700 verbs.
-The document does not specify whether these ‘‘verbs’’ are only verb roots or also other type of roots, such as nominals or adjectivals, nor how the parser works in detail.
+The document does not specify whether these “verbs” are only verb roots or also other type of roots, such as nominals or adjectivals, nor how the parser works in detail.
 According to some information that has been gathered, the parser would be developed using the Perl programming language and would also be able to generate sentences in Mapudüngun.
@@
 1.1.3   A Lexical Study
-In her thesis work (2016), Villena states that her research focuses on examining the authenticity and efficacy of the processes employed in Mapudüngun noun formation.
+In her thesis work Villena (2016) states that her research focuses on examining the authenticity and efficacy of the processes employed in Mapudüngun noun formation.
 Her objective is to develop criteria for the creation of new nouns in this language.
@@
 However, the abstract fails to mention that the noun formation under investigation is mainly by derivation processes (compounding is also taken into account), that is to say, by the mechanisms this language commonly uses for the purpose of identifying entities [^11] From our point of view, these are not strictly neologisms, but available (morpho-syntactic) strategies that this language has to deal with the naming of things, new or known.
-For example, it would not be considered an English neologism: ‘Internet surfer’, even though ‘Internet’ is already a neologism, but ‘Internaut’ would be; it is not registered but possible.
-.
+For example, it would not be considered an English neologism: ‘Internet surfer’, even though ‘Internet’ is already a neologism, but ‘Internaut’ would be; it is not registered but possible..
 In any case, the subject of neologicity filters is addressed further down in the work, with the assertion that one of the criteria for determining whether a form is neological is the internal structure of the form (ibid: 11).
 This suggests that the root component of the neologism should be accurately lexically classified.
-However, the prospect of a ‘‘roots classification process’’ appears to have been definitively lost after this note: ‘‘In order to confirm the appropriateness of the lexical categories assigned by Augusta, the most adequate procedure would have been to verify their functioning in discourse.
-However, the size of the corpus to be analysed made this impossible’’ (ibid: note 82, p.
+However, the prospect of a “roots classification process” appears to have been definitively lost after this note: “In order to confirm the appropriateness of the lexical categories assigned by Augusta, the most adequate procedure would have been to verify their functioning in discourse.
+However, the size of the corpus to be analysed made this impossible” (ibid: note 82, p.
 114).
-–In the following point §2.1 we explain the procedure followed in our work, which actually implies the process mentioned in Villena’s note–.
-Moreover, it appears that accurate lexical classification is not of paramount importance in this research study, particularly in the context of composition, as evidenced by the following citation: ‘‘This finding led to the conclusion that the categories of roots involved in compounding are not crucial in determining the structure of Mapudüngun compounds.
-Instead, the type of semantic condition maintained between them is deemed to be more significant’’ [^15] As Villena (2016) notes, this is a conclusion of Baker and Fasola (2009) that is supported by the observation of a similarity in the order of the V-N and N-N compounds.
+–In the following point §2.1 we explain the procedure followed in our work, which actually implies the process mentioned in Villena (2016)’s note–.
+Moreover, it appears that accurate lexical classification is not of paramount importance in this research study, particularly in the context of composition, as evidenced by the following citation: “This finding led to the conclusion that the categories of roots involved in compounding are not crucial in determining the structure of Mapudüngun compounds.
+Instead, the type of semantic condition maintained between them is deemed to be more significant” [^12] As Villena (2016) notes, this is a conclusion of Baker and Fasola (2009) that is supported by the observation of a similarity in the order of the V-N and N-N compounds.
 (ibid: 63).
@@
 However, ultimately, we were only able to find confirmation of a task that had not been realised, as evidenced by the following few examples:
-E.03) (Villena, 2016)
+E.03
+(Villena, 2016)
   1. a.
@@
     yewen       (‘shame’)          ‘male or female genitalia’
-Example E.03.a ((6) b., p.: 53 in the source) shows küdaw- as a verbal root being nominalised by the suffix -fe, which the author says that combines with verbal bases to generate nouns that denote a characteristic agent.
+Example E.03a ((6) b., p.: 53 in the source) shows küdaw- as a verbal root being nominalised by the suffix -fe, which the author says that combines with verbal bases to generate nouns that denote a characteristic agent.
 But in the dictionaries consulted on the CORLEXIM website (Chandía, 2014) the following entries are found:
-  1. küdaw s. The work. (Augusta, F. de (2016a) map-cas)
-  2. küdaw N. work. (Smeets, Ineke (2018) map-ing)
-  3. küdaw work. (Febrés, A. (1846a) map-cas)
-  4. absorto adj. [To be] ngoymaduamkülen (intr.); ngoymakonkülen kiñe küdaw mew. (Augusta, F. de (2016b) cas-map)
-As demonstrated above, three authors identify the form as a noun; de Augusta provides an example of the use of the word ‘absorto’ in Mapudüngun (being engrossed in a job), where küdaw is used as a noun without any suffixes, which is in contrast to the classification used by Villena.
+  1. 1.
+    küdaw s. The work. (de Augusta map-cas in Chandía, 2014)
+  2. 2.
+    küdaw N. work. (Smeets map-ing in Chandía, 2014)
+  3. 3.
+    küdaw work. (Febrés map-cas in Chandía, 2014)
+  4. 4.
+    (to be) engrossed adj. ngoymaduamkülen; (de Augusta cas-map in Chandía, 2014)
+    ngoymakonkülen kiñe küdaw mew ‘to be engrossed in a task".
+As demonstrated above, three authors identify the form as a noun; de Augusta provides an example of the use of the word ‘absorto’ in Mapudüngun (being engrossed in a job), where küdaw is used as a noun without any suffixes, which is in contrast to the classification used by Villena (2016).
 This example is one of many such instances that can be found in her study.
-E.03.b ((92) g., p.: 170 in the source) has been identified as an example of sintagmation, and it is said to represent the Prep-N-V structure. However, a review of dictionaries reveals that:
-  1.
+E.03b ((92) g., p.: 170 in the source) has been identified as an example of sintagmation, and it is said to represent the Prep-N-V structure. However, a review of dictionaries reveals that:
 pu- Vi.
@@
 puw- is most frequent.
-I found pu-fu-n ‘I arrived, but’ (RR) and iñche pu-n ‘I arrived’ (JM), and I found puw-küle- (RR) only once; müna ayuwüy iñchiu yu puwel ‘he was very glad that we had arrived’; küpa pulen liwen ‘I want to arrive early’; petu puwlaymi küdawmew ‘you are too young to work’; (Smeets, Ineke (2018) map-ing)
-Literally translated, E.03.b would mean something like: ‘arrived at the ground as he was going’ or ‘reached the ground on his way’.
+I found pu-fu-n ‘I arrived, but’ (RR) and iñche pu-n ‘I arrived’ (JM), and I found puw-küle- (RR) only once; müna ayuwüy iñchiu yu puwel ‘he was very glad that we had arrived’; küpa pulen liwen ‘I want to arrive early’; petu puwlaymi küdawmew ‘you are too young to work’ (Smeets map-ing in Chandía, 2014.)
+Literally translated, E.03b would mean something like: ‘arrived at the ground as he was going’ or ‘reached the ground on his way’.
 Perhaps there has been some confusion with another homophonic particle which is a prepositive and which de Augusta and Febrés, among others, define as follows:
-  1. pu- artic. pref. pluralises the names of people, animals and some things. pu wentru ‘the men’. pu wangül‘en ‘the stars’. See Augusta (1903, p. 15, 1). (Augusta, F. de (2016a) map-cas)
-  2.
-pu- (part.
+pu artic. pref. pluralises the names of people, animals and some things. pu wentru ‘the men’. pu wangülen ‘the stars’. See Augusta (1903, p. 15, 1) (de Augusta map-cas in Chandía, 2014.)
+pu (part.
 bef.
 nouns) denotes plurality; item, at or in; by; under; e.g.
-pu ovida, sheep (pl.); pu ruka, at home, in the house; pu liwen, in the morning; pu minue, under the ground.
-(Febrés, A.
-(1846a) map-cas)
-In addition, there is a high probability that the translocative suffix -pu- derives from the same verb, pu-.
+pu ofida, sheep (pl.); pu ruka, at home, in the house; pu liwen, in the morning; pu minue, under the ground (Febrés map-cas in Chandía, 2014)
+In addition, there is a high probability that the translocative suffix -pu derives from the same verb, -pu.
 The affix indicates that the denoted event takes place at a location away from the speaker.
 This function has also been described by Febrés –and others–:
-  1.
 pu, rpu (interp.
@@
 gr.
-pepuavin or perpuavin, I’ll drop by to see him; pipuavin or pirpuavin, I’ll pass on to say to him.
-(Febrés, A.
-(1846a) map-cas)
-E.03.c ((98) b., p.: 176 in the source) is employed to illustrate a case of metonymy, yet it transpires to be a case of semantic calque from the Spanish ’las vergüenzas’.
-In point of fact, the online DLE dictionary provides the eighth definition for the entry "vergüenza" as follows:
-  1.
+pepuafin or perpuafin, I’ll drop by to see him; pipuafin or pirpuafin, I’ll pass on to say to him (Febrés map-cas in Chandía, 2014)
+E.03c ((98) b., p.: 176 in the source) is employed to illustrate a case of metonymy, yet it transpires to be a case of semantic calque from the Spanish ‘las vergüenzas’.
+In point of fact, the online DLE dictionary provides the eighth definition for the entry “vergüenza” as follows:
 f.
@@
 While lexicons represent the most valuable resource in this respect, the remaining works also offer valuable insights that can facilitate the resolution of ambiguities in numerous cases.
+1.2   Summary of the Work
+This study is a systematic reclassification of the lexical categories of roots involved in Mapudüngun verbal themes.
+The primary goal of the work is to establish a reliable foundation for valency classification (Chandía, 2026) by ensuring that roots identified as verbal are indeed inherently verbal, rather than verbalised forms belonging to other lexical categories (e.g., nouns, adjectives, adverbs).
+The methodology employed is a straightforward verification process: roots listed as verbal in Smeets (2008)’s dictionary are cross-referenced against their usage in isolation in historical dictionaries (de Augusta, Febrés, de Valdivia) and the autobiography of the longko Paskwal Koña (Wilhelm et al., 1930).
+If a root is found used without derivational suffixes in these sources, it is inferred that its original category is non-verbal.
+This process yielded a comprehensive table (Table 3 in the 4   Appendix) showing 238 roots initially classified as verbs but reclassified as nouns, adjectives, adverbs, or demonstratives.
+Key findings from this study include:
+  • A significant number of roots in Smeets (2008)’s verbal inventory are verbalised forms of other lexical categories.
+  • Lists of verbal roots that co-occur with the causative suffix -(ü/e)l and -(ü)m were compiled from the longko Koña’s memoirs (Wilhelm et al., 1930).
+  • A set of labile roots (those that can be used both transitively and intransitively without morphological change) was identified.
+The following article (Chandía, 2026) builds directly on this foundation. The lexical classifications established here are taken as the starting point for the analysis that follows.
 2   Verbal Roots
-As mentioned in the Introduction1, p.
-1), for a computational system like the one developed here, correctly identifying the valency of verbal roots is essential for computational morphological analysis.
+As mentioned in the 1   Introduction1), for a computational system like the one developed here, correctly identifying the valency of verbal roots is essential for computational morphological analysis.
 Many of the rules governing the combination of roots and suffixes depend on this classification or on changes in value during the word formation process.
@@
 Typically, the sentence context resolves the ambiguity, but this task must be handled by another system: a grammatical disambiguator.
-However, this tool is only helpful when analysing sentences; for isolated words, or words without context, it is less useful.
-.
+However, this tool is only helpful when analysing sentences; for isolated words, or words without context, it is less useful..
 As stated in the introduction, before undertaking the task of valency classification for Mapudüngun verbal roots, a (re)categorisation of the roots that form verbal themes in this language will be carried out.
-The aim is to more accurately determine the lexical category of these roots, which will help in identifying the roots to be included in the subsequent work on recognising verbal valency (Chandía, 2025).
+The aim is to more accurately determine the lexical category of these roots, which will help in identifying the roots to be included in the subsequent work on recognising verbal valency (Chandía, 2026).
 From a computational perspective, this will also enable a better morphological analysis of Mapuche verbs.
 2.1   Lexical Category of Roots
-The verbal roots listed by Smeets in her dictionary were examined to verify their lexical classification and to modify it if they actually belong to other lexical categories and have been verbalised instead.
+The verbal roots listed by Smeets (2008) in her dictionary were examined to verify their lexical classification and to modify it if they actually belong to other lexical categories and have been verbalised instead.
 To achieve this, we primarily referred to the dictionaries by de Augusta, Febrés, and de Valdivia through the CORLEXIM (Chandía, 2014) platform, as well as to the autobiography of the longko [^15] Literally, longko means ‘head’, but similar to other languages, such as Catalan, it is used to denote a chief or leader.
@@
 In a more everyday sense, toki means ‘axe’.
-Paskwal Koña [^19] The longko Paskwal Koña narrated his life to the Capuchin missionary Ernst Wilhelm von Mösbach, who transcribed it in Mapudüngun.
+Paskwal Koña [^16] The longko Paskwal Koña narrated his life to the Capuchin missionary Alois Wilhelm von Mösbach (padre Ernesto), who transcribed it in Mapudüngun.
 This took place in the 1920s when Koña was still an energetic longko of 80 years.
@@
 If any root was found used in isolation —i.e., not as part of a verbal form, nor with affixes attached— it was inferred that the original form is nominal, adjectival, or adverbial, depending on the context and usage.
-Many of these roots also appear in Smeets’ dictionary with the category listed in the column ‘Final Cat.’ in Table LABEL:tab1.
+Many of these roots also appear in Smeets (2008)’s dictionary with the category listed in the column ‘Final Cat.’ in Table 1.
 These are non-verbal lexical categories, indicating that these words have been verbalised.
@@
 This column also reflects the lexical category reached after verifying each root in this study.
-The categories of the roots extracted from Smeets’ dictionary are listed in the ‘Initial Cat.’ column.
-Some of these roots are recorded by Smeets with an incorporated suffix, which can alter their meaning and valency when the suffix is absent.
-These suffixes are extracted from the original form and listed in the ‘Extracted suffixes’ column –the following Table LABEL:tab1 is only a sample available for easy verification of what is stated here.
-The complete table, which contains two hundred and thirty-eight roots, can be found in the Appendix, §4, Table LABEL:tab3, p.
-LABEL:tab3–.
-Table 1: Verbal Roots Reclassified As Other PoS (Sample)
-Root Extracted
-Suffixes
-Initial
-Cat.
-Final
-Cat.
-Definition
-ad- VI N the visible or characteristic aspects of something or someone, or of a group
-chadi- VI N salt
-dakel- VT N commission, agreement
-[…]
-fa- VI Dem this, here
-ina- VI Av close
-kachu- VI N grass, herb
-la- VI Aj dead, deceased
-llag- VT Aj half
-machi- VI N shaman
-nag- VI Av down, below, under, beneath
-ngeñi- -ka- VI Aj urgent, pressing
-ñochi- VI Aj peaceful, calm
-paf- VI N boil, abscess
-raküm- VT N closure, wall, protection
-tonon- VT N weft
-traf- VI Aj together, gathered, joined, fitted, interlocked
-[…]
-üna- VI N itch
-wachi- VT N trap
-yafü- VI Aj hard, firm
+The categories of the roots extracted from Smeets (2008)’s dictionary are listed in the ‘Initial Cat.’ column.
+Some of these roots are recorded by Smeets (2008) with an incorporated suffix, which can alter their meaning and valency when the suffix is absent.
+These suffixes are extracted from the original form and listed in the ‘Extracted suffixes’ column —the following Table 1 is only a sample available for easy verification of what is stated here.
+The complete table, which contains two hundred and thirty-eight roots, can be found in the 4   Appendix, §4, Table 3.
+?tablename? 1: Verbal Roots Reclassified as Other Parts of Speech (Sample)
+Root Extracted Initial Final Definition
+Suffixes Cat. Cat.
+ad- IV N the visible or characteristic aspects of something or someone, or of a group
+chadi- IV N salt
+dakel- TV N commission, agreement
+echid- $\sim$ echiw- IV N sneeze
+fa- IV Dem this, here
+ina- IV Av close
+kachu- IV N grass, herb
+la- IV Aj dead, deceased
+llag- TV Aj half
+machi- IV N shaman
+nag- IV Av down, below, under, beneath
+ngeñi- -ka IV Aj urgent, pressing
+ñochi- IV Aj peaceful, calm
+paf- IV N boil, abscess
+raküm- TV N closure, wall, protection
+tonon- TV N weft
+traf- IV Aj together, gathered, joined, fitted, interlocked
+uma- $\sim$ umañ- $\sim$ umag- $\sim$ umaw- IV N sleepiness, drowsiness
+üna- IV N itch
+wachi- TV N trap
+yafü- IV Aj hard, firm
+Abbreviations: IV = Intransitive Verb; TV = Transitive Verb; N = Noun; Aj = Adjective; Av = Adverb;
+Dem = Demonstrative.
 2.2   Valency of Verbal Roots
 The second step in the classification of roots that participate in Mapuche verbal themes is to specify the valency of verbal roots; as previously mentioned, this task will be presented in Chandía (2026).
-However, we will preview the results of this study and present a table of some verbal roots that co-occur with the causative suffix -(ü/e)l- (see the list of roots in p.
-2.3).
-We also provide a list of labile roots incorporated into Düngupeyüm and a table of roots that undergo causativisation with the causative suffix -(ü)m- (Table LABEL:tab4).
-2.3 Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l-
+However, we will preview the results of this study and present a list of some verbal roots that co-occur with the causative suffix -(ü/e)l (see §2.3).
+We also provide a list of labile roots (see §2.4) and a sample table of roots that undergo causativisation with the causative suffix -(ü)m (Table 2).
+2.3   Verbal Roots That Co-occur With The Causative Suffix -(ü/e)l
 In the autobiography of the longko (Wilhelm et al., 1930), twenty-one verbal roots [^17] Approximately one hundred and ten verbal roots that were not forming compounds were checked.
-Since this task took too long, it was decided to postpone the verification of roots when they participate in verbal compounds and exhibit a suffix in one of the forms that the causative can take: -l-, -ül-, -el-.
-were found that combine with the causative -l-, as shown in the list bellow.
-Almost all the roots are intransitive, except for two that form both transitive and intransitive verbs, meaning they are labile or bivalent roots (for a detailed explanation see §3.4 and §3.4.1 in Chandía, 2025).
-In the following list, the two senses of the labile verbs are marked according to their valency, with -IV (intransitive verb) and -TV (transitive verb).
+Since this task took too long, it was decided to postpone the verification of roots when they participate in verbal compounds and exhibit a suffix in one of the forms that the causative can take: -l, -ül, -el.
+were found that combine with the causative -l, as shown in the list below.
+Almost all the roots are intransitive, except for two that form both transitive and intransitive verbs, meaning they are labile or bivalent roots (for a detailed explanation see the companion valency article, Chandía, 2026).
+In the following list, the two senses of the labile verbs are marked according to their valency, with IV (intransitive verb) and TV (transitive verb).
   1. 1.
@@
   8. 8.
-    monge- ‘-IV to live, to heal, to get healed’ / ‘-TV to revive, to heal’
+    monge- IV ‘to live, to heal, to get healed’ / TV ‘to revive, to heal’
   9. 9.
@@
   10. 10.
-[…]
+    nepe- [^18] The computational system generates all possible analyses, which sometimes leads to conjectures about the origin of certain Mapudüngun forms.
+For instance, nepe- meaning ‘to wake up’ might have originated from the compound ne- $\sim$ nie- + pe: ‘to have’ + ‘sight, vision’.
+Similarly, it can be inferred that pewma meaning ‘dream’ comes from pe- + uma(g): ‘sight, vision’ + ‘sleep’.
 ‘to wake up’
@@
   20. 20.
-    wüño- ‘to return, to recover’
+    wüño- ‘to return, to start again’
   21. 21.
-    yewe- ‘-IV to feel embarrassed, to be ashamed’ / ‘-TV to respect’
+    yewe- IV ‘to feel embarrassed, to be ashamed’ / TV ‘to respect’
 2.4   Collecting Labile Roots In Düngupeyüm
-In the Düngupeyüm morphological analysis system, labile roots will now appear in both the transitive and intransitive lists, which inevitably leads to increase ambiguity when analysing verb forms; however, excluding these roots would omit potential analyses.
-Additionally, the system has been updated to ensure that causatives, whether -(ü)m- or -(ü)l- (Chandía, 2008, §3.3 and §3.4 respectively), do not co-occur with transitive roots.
+In the Düngupeyüm morphological analysis system, labile roots will now appear in both the transitive and intransitive lists, which inevitably leads to increased ambiguity when analysing verb forms; however, excluding these roots would omit potential analyses.
+Additionally, the system has been updated to ensure that causatives, whether -(ü)m or -(ü)l, do not co-occur with transitive roots.
 Thus, causatives will only combine with labile roots in their intransitive sense.
@@
 List of labile roots collected in the system
-  • aye- ‘to laugh’ (-IV)
-  • aye- ‘to laugh at; to laugh with’ (-TV)
-  • kewa- ‘to fight, to quarrel, to struggle, to combat’ (-IV)
-  • kewa- ‘to hit, to strike, to defeat, to overcome’ (-TV)
-  • llüka- ‘to get frightened, to be scared, to fear, to be afraid’ (-IV)
-  • llüka- ‘to scare, to frighten’ (-TV)
-  • meke- ‘to take care of, to deal with, to be busy, to last, to endure’ (-IV)
-  • meke- ‘to make’ (-TV)
-  • monge- ‘to get life, to heal, to recover’ (-IV)
-  • monge- ‘to revive, to bring back to life, to heal’ (-TV)
-  • nge- ‘to be’ (-IV)
-  • nge- ‘to have’ (-TV) (see nge- -IV ‘to be’, -TV ‘to have’ in Chandía, 2025)
-  • püna- ‘to stick, to glue, to adhere’ (-IV)
-  • püna- ‘to glue, to paste’ (-TV)
-  • waychüf- ‘to fall, to get turned over, to roll, to spin’ (-IV)
-  • waychüf- ‘to discard, to flip, to roll, to spin’ (-TV)
-  • yewe- ‘to feel embarrassed, to be ashamed’ (-IV)
-  • yewe- ‘to respect’ (-TV)
-2.5 Verbal Roots That Co-occur With The Causative Suffix -(ü)m-
-In Chandía, 2025: §3.3, it will be shown that verifying the presence of the causative suffix -(ü)m- in Mapuche verb forms is a reliable method for determining that the root carrying it is intransitive.
-However, it is also necessary to verify whether the root is genuinely verbal or it belongs to another lexical category, a task that this document addresses.
-Furthermore, as noted by Longkon (2011a), Golluscio (2007) and Smeets, not all intransitive roots accept causativisation with -(ü)m-.
+  1. 1.
+    aye- ‘to laugh’ (IV)
+    aye- ‘to laugh at; to laugh with’ (TV)
+  2. 2.
+    kewa- ‘to fight, to quarrel, to struggle, to combat’ (IV)
+    kewa- ‘to hit, to strike, to defeat, to overcome’ (TV)
+  3. 3.
+    llüka- ‘to get frightened, to be scared, to fear, to be afraid’ (IV)
+    llüka- ‘to scare, to frighten’ (TV)
+  4. 4.
+    meke- ‘to take care of, to deal with, to be busy, to last, to endure’ (IV)
+    meke- ‘to make’ (TV)
+  5. 5.
+    monge- ‘to get life, to heal, to recover’ (IV)
+    monge- ‘to revive, to bring back to life, to heal’ (TV)
+  6. 6.
+    nge- ‘to be’ (IV)
+    nge- ‘to have’ (TV) (see §3.4.2.3 in Chandía, 2026)
+  7. 7.
+    püna- ‘to stick, to glue, to adhere’ (IV)
+    püna- ‘to glue, to paste’ (TV)
+  8. 8.
+    waychüf- ‘to fall, to get turned over, to roll, to spin’ (IV)
+    waychüf- ‘to discard, to flip, to roll, to spin’ (TV)
+  9. 9.
+    yewe- ‘to feel embarrassed, to be ashamed’ (IV)
+    yewe- ‘to respect’ (TV)
+2.5   Verbal Roots That Co-occur With The Causative Suffix -(ü)m
+In the companion valency article (Chandía, 2026): §3.3), it will be shown that verifying the presence of the causative suffix -(ü)m in Mapuche verb forms is a reliable method for determining that the root carrying it is intransitive.
+However, it is also necessary to verify whether the root is genuinely verbal or belongs to another lexical category, a task that this document addresses.
+Furthermore, as noted by Longkon (2011), Golluscio (2007) and Smeets (2008), not all intransitive roots accept causativisation with -(ü)m.
 Therefore, it is important to examine these authors’ claims regarding the role of telicity or atelicity in the selection of the causative suffix.
-In Table LABEL:tab4, all the roots that have been verified to combine with the causative suffix -(ü)m- are presented, the following Table LABEL:tab2 is only a sample available for easy verification of what is stated in Chandía, 2008: §3.3.
-Table 2: Roots That Causativise With -(ü)m- (Sample)
+In Table 4 (4   Appendix), all the roots that have been verified to combine with the causative suffix -(ü)m are presented.
+The following Table 2 is only a sample available for easy verification of what is stated in the valency article.
+?tablename? 2: Roots That Causativise With -(ü)m (Sample)
 Root Cat. Intransitive sense Root+CA Transitive sense Collected froma
-ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
-anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
-fa Dem this, here fa-m- to be like this K. S.
-firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
-kon VI to enter, to begin kon-üm- to make come in, to add K. G.
-la Aj dead, deceased lang-üm- to kill K. S. G.
-(l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
-llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
-[…]
-nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
-pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
-pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
-puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
-trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
-tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
-üy N name üy-üm- to name, to denominate S.
-wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
-yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
+ad N form, place, familiar, face, image, custom ad-üm to be experienced/skilled (in something) K. S.
+anü IV to get settled, to get sit anü-m to settle, to set, to put, to place K. S. G.
+fa Dem this, here fa-m to be like this K. S.
+firkü Aj fresh, cool firkü-m to cool, to warm slightly, to take the chill off K.
+kon IV to enter, to begin kon-üm to make come in, to add K. G.
+la Aj dead, deceased lang-üm to kill K. S. G.
+(l)el IV to let oneself, to become loose, to detach (l)el-üm to release, to let go, to kill K.
+llum IV to get hidden, to conceal oneself llum-üm to hide, to conceal K.
+nel Aj loose, untied, unleashed nel-üm $\sim$ nelk-üm to release, to let go, to free S.
+nor Aj straight, correct, right nor-üm to correct, to rectify, to straighten S.
+pewü IV to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m to measure, to reach out using the arms, to twist, to turn K.
+pümü Aj tense, tight, taut, strained pümü-m to stretch, to tighten, to tense K.
+puw IV to arrive, to reach, to achieve, to stay, to remain puw-üm to send, to deliver, to convey, to serve K.
+trar IV to ooze, to suppurate, to secrete trar-üm to drain the pus S.
+tüng IV to calm down, to take time tüng-üm to leave alone, not to disturb K.
+üy N name üy-üm to name, to denominate S.
+wim Aj accustomed, used to, meek, tame, domesticated wim-üm to make accustomed, to domesticate, to tame S.
+yung Aj pointed, sharp yung-üm to sharpen, to make pointed K.
+aThe “Collected from” column indicates which source contains examples of the root with the causative suffix.
+K. = Wilhelm et al., 1930 (autobiography of longko Paskwal Koña); S. = Smeets, 2008 (A Grammar of Mapuche); G. = Golluscio, 2007 (Gramática descriptiva del mapuche).
 3   Conclusions
-In the present work, we undertook a reclassification of the lexical categories of roots involved in verbal themes of the Mapuche language, which are part of the inventory of examples in the descriptive grammar of Mapudüngun developed by the anthropologist Ineke Smeets.
-This is a preliminary task to valency classification (Chandía, 2025), the study that motivated us to begin this lexical category review.
-The task of valency classification of Mapuche verbal roots required an accurate classification of the lecixal categories of the roots to avoid attributing verbal features, such as valency, to elements that cannot exhibit them because they are not verbal.
-A simple verification method for the lexical category was devised, as explained in §2.1, p.2.1, and despite its simplicity, it has proven to be effective, with results shown in Table LABEL:tab3, p.
-LABEL:tab3).
+In the present work, we undertook a reclassification of the lexical categories of roots involved in verbal themes of the Mapuche language, which are part of the inventory of examples in the descriptive grammar of Mapudüngun developed by the anthropologist Ineke Smeets (2008).
+This is a preliminary task to valency classification (Chandía, 2026), the study that motivated us to begin this lexical category review.
+The task of valency classification of Mapuche verbal roots required an accurate classification of the lexical categories of the roots to avoid attributing verbal features, such as valency, to elements that cannot exhibit them because they are not verbal.
+A simple verification method for the lexical category was devised, as explained in §2.1, and despite its simplicity, it has proven to be effective, with results shown in Table 3 (4   Appendix).
 Next, as the two studies referred to in this article are now complete, we present some of the overall material here, as we have included some of the data generated on completion of both studies in this initial document.
 In a computational system like an automatic morphological analyser, where it is crucial to have specific and non-redundant data, ambiguity can be significantly reduced if roots are listed, as much as possible, under a single classification per form.
-Despite this, we chose to list roots with ambivalent behaviour (see §2.4, p.
-2.4) in both lists, intransitive verbal roots and transitive verbal roots, with each list reflecting the root’s valency-related sense.
+Despite this, we chose to list roots with ambivalent behaviour (see §2.4) in both lists, intransitive verbal roots and transitive verbal roots, with each list reflecting the root’s valency-related sense.
 This approach might introduce varying degrees of ambiguity in the analyses produced by the machine.
 However, besides the ambiguity, it also allows for potential analyses that would be excluded if ambivalent roots were not collected in this manner.
-It is essential to first identify the category most likely to be the source of the multiple interpretations or derivations that a form, used as the main component of the verbal theme, might generate (see §2.1, p.
-2.1).
-Therefore, the primary task is to ensure that valency values are sought in inherently verbal roots rather than in verbalised roots derived from other categories (see Table LABEL:tab3, p.
-LABEL:tab3).
+It is essential to first identify the category most likely to be the source of the multiple interpretations or derivations that a form, used as the main component of the verbal theme, might generate (see §2.1).
+Therefore, the primary task is to ensure that valency values are sought in inherently verbal roots rather than in verbalised roots derived from other categories (see Table 3 in the 4   Appendix).
 This can be challenging in a language like Mapudüngun, as there is often insufficient material or stored knowledge to determine the etymology of much of its vocabulary, largely due to historical and social factors.
@@
 Nonetheless, it is at least known that Wilhelm et al.
-(1930) is one of the best references for the Mapuche language (see footnote 19, p.
-19), which was the main reason for its use in this research.
-To conclude, the decision to publish the study in two parts —lexical classification presented here and valency classification presented in Chandía, 2025— aims to clarify the findings.
-As stated in the summary, ‘‘The outcomes of this categorical and valency review directly benefit the computational analyser, as they are implemented as soon as they are verified.
-Additionally, these findings are expected to enhance the understanding of the Mapuche language’’.
+(1930) is one of the best references for the Mapuche language (see footnote 16), which was the main reason for its use in this research.
+To conclude, the decision to publish the study in two parts —lexical classification presented here and valency classification presented in Chandía, 2026— aims to clarify the findings.
+As stated in the summary, “The outcomes of this categorical and valency review directly benefit the computational analyser, as they are implemented as soon as they are verified.
+Additionally, these findings are expected to enhance the understanding of the Mapuche language”.
 Readers are invited to review the accompanying article to this first work, which provides a comprehensive understanding of the results of this research.
@@
 +2 Second person
-+3   Third person
++ Third person
 +3A Third person agent
@@
 +NU Numeral
-+$O$   Oblique object
++OO Oblique object
 +OVN Objective verbal noun
@@
 +TR Transitiviser
-VI, -IV   Intransitive verb
+IV, -IV Intransitive verb
 +VRB Verbaliser
-VT, -TV   Transitive verb
+TV, -TV Transitive verb
 4.2   Tables
-Table 3: Verbal Roots Reclassified As Other PoS
-Root Extracted
-Suffixes
-Initial
-Cat.
-Final
-Cat.
-Definition
-ad- VI N the visible or characteristic aspects of something or someone, or of a group
-af- VI N end
-afü- VI Aj cooked, ripe
-allush- VI Aj lukewarm
-alü- VI Av very, much
-angim- VT N jerky
-angka- VI N middle, half
-antü- VI N Sun, day, time
-are- VI Aj hot, warm
-ariñ- VI Aj burnt, smoked
-arof- VI N sweat
-asul- VI Aj blue
-awi- VI Aj hot
-awka- VI Aj rebellious, wild
-aylen- VI N ember(s)
-aywiñ- VI N shade
-chadi- VI N salt
-chafo- VI N cough
-chali- VI N greeting
-challwa- VI N fish
-chape- VT N braid (in the hair)
-che- VI N person, people
-chedkuy- VT N father/son in law (of a man)
-chilla- VT N saddle, chair
-chod- VI Aj yellow
-chum- VT Int how, what
-chüngküd- VI Aj circular, rounded
-dakel- VT N commission, agreement
-dew- VI Av already, since, because, no more, only, finally
-domo- VI N woman
-[…]
-duam- VT N matter, issue, need, intention
-dumdum- VI N penumbra
-düngu- VT N issue, matter, subject, concern, word, language
-[…]
-elfal- VT N commission
-eñum- VI Aj comfortingly/pleasantly warm
-epew- VI N tale, story, history, narration
-fa- VI Dem this, here
-fe- VI Dem that, there
-fente- VT Av so, much
-fentre- VI Av very, much
-fey- VI Dem this, that, she, he
-filla- VI N scarcity, lack
-fitrun- VI N smoke
-fücha- VI Aj big, old
-fülang- VI Aj white
-funa- VI Aj rotten
-füre- VI Aj spicy, hot
-futrül- VT N heap, pile
-füw- VT N thread, wool
-ina- VI Av close
-ingka- VT N defender
-kachu- VI N grass, herb
-kafe- VI N coffee
-kallfü- VI Aj blue, purple, violet
-kangka- VT Aj roasted
-karü- VI Aj green
-kashü- VI Aj gray
-katrü- VT Aj cut off, interrupted
-kawe- VI N oar
-kellu- VT N help
-kelü- VI Aj red
-kim- VT Aj wise, knower
-kochü- VI Aj sweet
-kofke- VT N bread
-kolü- VI Aj rust, reddish-brown (colour)
-kompañ- VT N partner
-koñoll- VI Aj purple, violet
-koñü- VI N baby, newborn
-korü- VI N soup, broth
-kotrü- VI Aj salty, sour, bitter
-kowkow- VI N owl
-küdaw- VI N work
-kude- VI N bet
-[…]
-küllay- -tu- VT N quillaya, quillay soap
-külü- VI Aj tilted, slanted
-küme- VI Aj good
-kunaw- VI Aj inflated, floating
-kupül- VT Aj fixed, secure
-kuram-b VI N egg
-küri- VI Aj lustful, lecherous
-kurü- VI Aj black
-kürüf- VI N wind
-kushe- VI Aj elderly, old (woman)
-kutran- VI N illness, disease, pain, suffering
-kütrüng- VT N bundle, bunch
-küyen- VI N moon, month
-la- VI Aj dead, deceased
-lantra- VI Aj bulging, thick
-lawen- VT N medicine (medicinal herb)
-lef- VI Aj fast (going/moving)
-leg- VI Aj certain, accurate, correct
-lichi- VI N milk
-lif- VI Aj clean
-[…]
-llag- VT Aj half
-llekü- VI Av close
-llitu- VI Aj initial, first
-[…]
-llüngüd- VT Aj flat, level
-llüpañ- VT Aj lying, reclining
-lüf- VI Aj scorched, roasted
-lüp-d -üm- VT Aj scorched, roasted
-lüykü- VI N drop
-machi- VI N shaman
-[…]
-mafü- VT N dowry
-[…]
-mangkü- VT N kick, boot
-mañum- VT N gratitude, thankfulness
-mawün- VT N rain
-may- VT Conj yes
-me- VI N excrement, scum, dross
-melkay- VI Aj slippery
-[…]
-minggako- VI N communal/community work
-misha- VT N dining companion
-[…]
-motri- VI Aj fat
-moy- VI N boil
-moyo- VT N breast
-muday- VI N chicha
-[…]
-müna- -le- VI Av many, much
-mupi- VI Aj true, genuine
-müpü- VI N wing
-[…]
-mütrüm- VT N call
-nag- VI Av down, below, under, beneath
-[…]
-ngeñi- -ka- VI Aj urgent, pressing
-ngoyma- VT Aj unproductive, neglected
-ngüdi- VT N obstruction
-ngülam- VT N advice, tip
-ngüñü- VI N hunger
-nor- VI Aj straight (in a physical or figurative sense), correct
-notu- -ka- VT N obstinacy, blindness
-nüla- VT Aj open
-nümü- VI N smell, odour
-[…]
-nürüf- VT Aj closed
-nütram- VT N recount, conversation, story
-nüyün- VI N tremor, quake
-ñochi- VI Aj peaceful, calm
-ñum- VI N upper millstone
-paf- VI N boil, abscess
-[…]
-pali- VI N paline ball
-panü- VT N fathom
-papel- VT N paper
-[…]
-perimo- -ntu- VI N omen
-perkü- VI N fart
-pewma- VI N daydream, dream
-pichi- $similar-to\sim püchü- simi$$similar-to\sim püt$ü- VI Aj little
-pidku- VT Aj stew (referring to a dish typically involving legumes cooked in a pot)
-piku- VI N North
-pilu- VI Aj deaf
-pire- VI N snow, hail
-[…]
-pitru- VI N scabies
-pod- VI Aj dirty
-pofre- VI Aj poor
-poy- VI N tumour, swelling
-püd- VI Aj thick, dense, condensed
-pülle- VI Av close, next
-püntü- VI N piece, chunk
-püñeñ- VI N daughter or son (to the mother or her sister)
-püro- VT N knot
-pütre- VT Aj blazing, scorching, burning, hot
-pütrü- VI Av in abundance, a lot, plentifully
-raküm- VT N closure, wall, protection
-rangiñ- VI N middle, half
-re- -le- VI Aj alone, exclusive, pure
-riku- VI Aj rich (it may be related to rükü- (further down)
-ruka- VT N house
-rükü- VI Aj mean, petty, greedy, miserly, tight-fisted, stingy
-rümpel- VI Aj jealous
-[…]
-rutra- -tu- VT N pinch
-tonon- VT N weft
-traf- VI Aj together, gathered, joined, fitted, interlocked
-traf-iya-f VI Av last night, since last night’s fall
-tralka- VI N thunder, shotgun
-tran- VI Aj fallen (especially referring to trees)
-trangliñ- VI N frost
-trapel- VT N cord, string
-trar- VI N pus
-[…]
-traytray- VI N gargle, or similar gargling sound
-trem- VI Aj grown, adult, mature
-tremo- VT Aj healthy, robust, splendid
-trintri- VI Aj curly
-tripantu- VI N year
-tromü- VI N cloud
-tror- VI N foam
-trüfon- VI N cough
-trufür- VI N dust
-[…]
-trülke- VT N leather
-[…]
-[…]
-trutruka- VI N a wind instrument like a long cornet
-tute- VT Aj correct, accurate
-üküm- VI N silence
-ül- VI N song, singing
-[…]
-[…]
-üna- VI N itch
-üng- -üm- VT AV while
-[…]
-[…]
-üyag- -tu- VT Aj both
-wachi- VT N trap
-waw- VI N valley, lowland, low-lying area
-wechod- VI N hole
-wedwed- VI Aj mischievous
-welu- VI Aj alternative, alternating
-wentru- VI N man
-wew- VT N victory, triumph, gain, profit
-[…]
-[…]
-wif- VI Aj wide, broad
-[…]
-wilüf- VI Aj shiny
-wim- VI Aj tame, accustomed
-wima- VT N rod, staff
-witral- VT N warp, loom
-wün- VI N mouth, dawn, sunrise
-wüne- VI Aj fore, anterior, superior
-wütre- VI Aj cold
-yafü- VI Aj hard, firm
-yall- VT N son
-[…]
-yochi- VI Aj tight, fitted
-yung- -üm- VT Aj pointed, sharp
-a. Küllay is a tree endemic to the central region of Chile; scientific name Quillaja saponaria, also known as ‘soapbark tree’.
-b.
-In conjugated (verbalised) forms, the root kuram ‘egg’ is alternatively analysed as -NN.kürastone+CA.m34… by Düngupeyüm, that is to say, as kura-m, a form that may be interpreted as ‘caused to be a stone’, which may well be the origin of kuram ‘egg’.
-[…]
-d.
-The causative suffix -(ü)m- triggers a radical change in some roots, causing those ending in f to shift to p (see Chandía, 2025, E.06), indicating that this root (lüp-) is the same as the one above (lüf-).
-e. Palin is a sport similar to field hockey, referred to as chueca by Spaniards and Creoles.
-[…]
-Table 4: Roots That Causativise With -(ü)m-
+?tablename? 3: Verbal Roots Reclassified as Other Parts of Speech
+Root Extracted Initial Final Definition
+Suffixes Cat. Cat.
+ad- IV N the visible or characteristic aspects of something or someone, or of a group
+af- IV N end
+afü- IV Aj cooked, ripe
+allush- IV Aj lukewarm
+alü- IV Av very, much
+angim- TV N jerky
+angka- IV N middle, half
+antü- IV N Sun, day, time
+are- IV Aj hot, warm
+ariñ- IV Aj burnt, smoked
+arof- IV N sweat
+asul- IV Aj blue
+awi- IV Aj hot
+awka- IV Aj rebellious, wild
+aylen- IV N ember(s)
+aywiñ- IV N shade
+chadi- IV N salt
+chafo- IV N cough
+chali- IV N greeting
+challwa- IV N fish
+chape- TV N braid (in the hair)
+che- IV N person, people
+chedkuy- TV N father/son in law (of a man)
+chilla- TV N saddle, chair
+chod- IV Aj yellow
+chum- TV Int how, what
+chüngküd- IV Aj circular, rounded
+dakel- TV N commission, agreement
+dew- IV Av already, since, because, no more, only, finally
+domo- IV N woman
+doy- $\sim$ yod- TV Av more
+duam- TV N matter, issue, need, intention
+dumdum- IV N penumbra
+düngu- TV N issue, matter, subject, concern, word, language
+echid- $\sim$ echiw- IV N sneeze
+elfal- TV N commission
+eñum- IV Aj comfortingly/pleasantly warm
+epew- IV N tale, story, history, narration
+fa- IV Dem this, here
+fe- IV Dem that, there
+fente- TV Av so, much
+fentre- IV Av very, much
+fey- IV Dem this, that, she, he
+filla- IV N scarcity, lack
+fitrun- IV N smoke
+fücha- IV Aj big, old
+fülang- IV Aj white
+funa- IV Aj rotten
+füre- IV Aj spicy, hot
+futrül- TV N heap, pile
+füw- TV N thread, wool
+ina- IV Av close
+ingka- TV N defender
+kachu- IV N grass, herb
+kafe- IV N coffee
+kallfü- IV Aj blue, purple, violet
+kangka- TV Aj roasted
+karü- IV Aj green
+kashü- IV Aj gray
+katrü- TV Aj cut off, interrupted
+kawe- IV N oar
+kellu- TV N help
+kelü- IV Aj red
+kim- TV Aj wise, knower
+kochü- IV Aj sweet
+kofke- TV N bread
+kolü- IV Aj rust, reddish-brown (colour)
+kompañ- TV N partner
+koñoll- IV Aj purple, violet
+koñü- IV N baby, newborn
+korü- IV N soup, broth
+kotrü- IV Aj salty, sour, bitter
+kowkow- IV N owl
+küdaw- IV N work
+kude- IV N bet
+kufü- $\sim$ kufüll- VI/TV Aj spoiled, damaged (due to the effect of heat)
+küllay- -tu TV N quillaya, quillay soap
+külü- IV Aj tilted, slanted
+küme- IV Aj good
+kunaw- IV Aj inflated, floating
+kupül- TV Aj fixed, secure
+kuram-b IV N egg
+küri- IV Aj lustful, lecherous
+kurü- IV Aj black
+kürüf- IV N wind
+kushe- IV Aj elderly, old (woman)
+kutran- IV N illness, disease, pain, suffering
+kütrüng- TV N bundle, bunch
+küyen- IV N moon, month
+la- IV Aj dead, deceased
+lantra- IV Aj bulging, thick
+lawen- TV N medicine (medicinal herb)
+lef- IV Aj fast (going/moving)
+leg- IV Aj certain, accurate, correct
+lichi- IV N milk
+lif- IV Aj clean
+ling-ar $\sim$ lüng-af $\sim$ lüg- $\sim$ lig- -ar, -afc IV Aj white, clear
+llag- TV Aj half
+llekü- IV Av close
+llitu- IV Aj initial, first
+lliwa- $\sim$ lluwa- -tu TV N soothsayer, seer
+llüngüd- TV Aj flat, level
+llüpañ- TV Aj lying, reclining
+lüf- IV Aj scorched, roasted
+lüp-d -üm TV Aj scorched, roasted
+lüykü- IV N drop
+machi- IV N shaman
+madom- $\sim$ masom- tuku- TV N ember, charred log
+mafü- TV N dowry
+mallma- $\sim$ malma- IV Aj boastful, arrogant, pretentious
+mangkü- TV N kick, boot
+mañum- TV N gratitude, thankfulness
+mawün- TV N rain
+may- TV Conj yes
+me- IV N excrement, scum, dross
+melkay- IV Aj slippery
+meñku- $\sim$ menku- TV N load, burden
+minggako- IV N communal/community work
+misha- TV N dining companion
+mollfü- $\sim$ mollfüñ- IV N blood
+motri- IV Aj fat
+moy- IV N boil
+moyo- TV N breast
+muday- IV N chicha
+mukur- $\sim$ mukud- IV Aj sour, bitter
+müna- -le IV Av many, much
+mupi- IV Aj true, genuine
+müpü- IV N wing
+mürke- $\sim$ murke- TV N toasted flour
+mütrüm- TV N call
+nag- IV Av down, below, under, beneath
+nel- $\sim$ ney- IV Aj loose, unfastened
+ngeñi- -ka IV Aj urgent, pressing
+ngoyma- TV Aj unproductive, neglected
+ngüdi- TV N obstruction
+ngülam- TV N advice, tip
+ngüñü- IV N hunger
+nor- IV Aj straight (in a physical or figurative sense), correct
+notu- -ka TV N obstinacy, blindness
+nüla- TV Aj open
+nümü- IV N smell, odour
+nüngay- $\sim$ ningay- IV N anger, fury, rage, impatience
+nürüf- TV Aj closed
+nütram- TV N recount, conversation, story
+nüyün- IV N tremor, quake
+ñochi- IV Aj peaceful, calm
+ñum- IV N upper millstone
+paf- IV N boil, abscess
+pakar- $\sim$ pakaka- IV N a large toad
+pali- IV N paline ball
+panü- TV N fathom
+papel- TV N paper
+payun- $\sim$ payuin- $\sim$ payum- $\sim$ payuñ- $\sim$ payun’ IV N beard
+perimo- -ntu IV N omen
+perkü- IV N fart
+pewma- IV N daydream, dream
+pichi- $\sim$ püchi- $\sim$ püchü- $\sim$ püti- $\sim$ pütü- IV Aj little
+pidku- TV Aj stew (referring to a dish typically involving legumes cooked in a pot)
+piku- IV N North
+pilu- IV Aj deaf
+pire- IV N snow, hail
+pitra- $\sim$ pitraw- IV N blister, callus
+pitru- IV N scabies
+pod- IV Aj dirty
+pofre- IV Aj poor
+poy- IV N tumour, swelling
+püd- IV Aj thick, dense, condensed
+pülle- IV Av close, next
+püntü- IV N piece, chunk
+püñeñ- IV N daughter or son (to the mother or her sister)
+püro- TV N knot
+pütre- TV Aj blazing, scorching, burning, hot
+pütrü- IV Av in abundance, a lot, plentifully
+raküm- TV N closure, wall, protection
+rangiñ- IV N middle, half
+re- -le IV Aj alone, exclusive, pure
+riku- IV Aj rich (it may be related to rükü- further down)
+ruka- TV N house
+rükü- IV Aj mean, petty, greedy, miserly, tight-fisted, stingy
+rümpel- IV Aj jealous
+runa- $\sim$ truna- TV N two-handed handful
+rutra- -tu TV N pinch
+tonon- TV N weft
+traf- IV Aj together, gathered, joined, fitted, interlocked
+traf-iya-f IV Av last night, since last night’s fall
+tralka- IV N thunder, shotgun
+tran- IV Aj fallen (especially referring to trees)
+trangliñ- IV N frost
+trapel- TV N cord, string
+trar- IV N pus
+trari- $\sim$ trarü- TV N ribbon, tape, belt
+traytray- IV N gargle, or similar gargling sound
+trem- IV Aj grown, adult, mature
+tremo- TV Aj healthy, robust, splendid
+trintri- IV Aj curly
+tripantu- IV N year
+tromü- IV N cloud
+tror- IV N foam
+trüfon- IV N cough
+trufür- IV N dust
+trüg- $\sim$ trow- $\sim$ trig- IV N crack
+trülke- TV N leather
+trümfül- $\sim$ trünfül- $\sim$ trüfül- IV Aj twisted, bent
+trür- $\sim$ mür- $\sim$ ür- IV N pair, couple
+trutruka- IV N a wind instrument like a long cornet
+tute- TV Aj correct, accurate
+üküm- IV N silence
+ül- IV N song, singing
+üllüf- $\sim$ illüf- IV Aj unfortunate, miserable, eventful
+uma- $\sim$ umañ- $\sim$ umag- $\sim$ umaw- IV N sleepiness, drowsiness
+üna- IV N itch
+üng- -üm TV Av while
+ütrüf- $\sim$ utruf- $\sim$ ütüf- $\sim$ itrüf- TV N throw (in a game)
+üy- $\sim$ güy- IV N name
+üyag- -tu TV Aj both
+wachi- TV N trap
+waw- IV N valley, lowland, low-lying area
+wechod- IV N hole
+wedwed- IV Aj mischievous
+welu- IV Aj alternative, alternating
+wentru- IV N man
+wew- TV N victory, triumph, gain, profit
+wichü- $\sim$ wüchü- $\sim$ wüchür- TV Aj twisted, crooked
+widi- $\sim$ widü- $\sim$ wüdü- TV N clay vessel/container, pottery
+wif- IV Aj wide, broad
+wikef- $\sim$ wikür- $\sim$ wünfü- $\sim$ wingür- IV/TV Aj torn, broken
+wilüf- IV Aj shiny
+wim- IV Aj tame, accustomed
+wima- TV N rod, staff
+witral- TV N warp, loom
+wün- IV N mouth, dawn, sunrise
+wüne- IV Aj fore, anterior, superior
+wütre- IV Aj cold
+yafü- IV Aj hard, firm
+yall- TV N son
+yiwül- $\sim$ yiwüll- TV N shuttle (referring to the loom)
+yochi- IV Aj tight, fitted
+yung- -üm TV Aj pointed, sharp
+a Küllay is a tree endemic to the central region of Chile; scientific name Quillaja saponaria, also known as ‘soapbark tree’.
+b In conjugated (verbalised) forms, the root kuram ‘egg’ is alternatively analysed as -nn.küra_stone+ca.m34… by Düngupeyüm, that is to say, as kura-m, a form that may be interpreted as ‘caused to be a stone’, which may well be the origin of kuram ‘egg’.
+c In this case (ling-ar-, lüng-af-), the suffixes were not extracted but are simply presented here.
+They appear to be suffixes (-ar-, -af-) that imply a semantic change (direction) when added to the base form.
+They might be unproductive suffixes that have become lexicalised with these roots, or they could be mere phonological fillers.
+No references to these elements were found in the various studies on Mapudüngun.
+However, it should be noted that a radical change occurs in the root lüg- $\sim$ lig-, similar in nature to that explained in the companion valency article (Chandía, 2026, E.06), though this is triggered by a suffix about which we have no information.
+d The causative suffix -(ü)m triggers a radical change in some roots, causing those ending in f to shift to p (see Chandía, 2026, E.06), indicating that this root (lüp-) is the same as lüf- above.
+e Palin is a sport similar to field hockey, referred to as chueca by Spaniards and Creoles.
+f trafiya $\sim$ trafüya is a lexicalised compound form, likely derived from traf- ‘to come together’ and wiya $\sim$ wüya ‘yesterday’, meaning ‘the coming together of yesterday [with today]’, referring to the period from the previous evening to the following morning, ‘last night’.
+?tablename? 4: Roots That Causativise With -(ü)m
 Root Cat. Intransitive sense Root+CA Transitive sense Collected froma
-1. ad N form, place, familiar, face, image, custom ad-üm- to be experienced/skilled (in something) K. S.
-2. afü Aj cooked, ripe afü-m- to cook K. S. G.
-3. ali VI to get heated up ali-m- to heat up S.
-4. angkü VI to get dry (up), to be sterile, to be arid angkü-m- to dry (up) K. S.
-5. anü VI to get settled, to get sit anü-m- to settle, to set, to put, to place K. S. G.
-6. are Aj hot, warm are-m- to warm up S. G.
-7. chaw N father, man chaw-üm-b to incubate K. S.
-8. che N person, people, human being che-m- to relatec K.
-9. chong VI to go out, to extinguish, to dry up chong-üm- to turn off, to extinguish, to dry S.
+1. ad N form, place, familiar, face, image, custom ad-üm to be experienced/skilled (in something) K. S.
+2. afü Aj cooked, ripe afü-m to cook K. S. G.
+3. ali IV to get heated up ali-m to heat up S.
+4. angkü IV to get dry (up), to be sterile, to be arid angkü-m to dry (up) K. S.
+5. anü IV to get settled, to get sit anü-m to settle, to set, to put, to place K. S. G.
+6. are Aj hot, warm are-m to warm up S. G.
+7. chaw N father, man chaw-ümb to incubate K. S.
+8. che N person, people, human being che-m to relatec K.
+9. chong IV to go out, to extinguish, to dry up chong-üm to turn off, to extinguish, to dry S.
 10. chu Int interrogative particle chu-m to ask a questiond K. S.
-11. fa Dem this, here fa-m- to be like this K. S.
-12. fay VI to ferment fay-üm- to [make] ferment S.
-13. fe Dem that, there fe-m- to be like that K. S.
-14. fey VI to get adapted, to fit fey-üm- to adapt, to fit, to insert S.
-15. firkü Aj fresh, cool firkü-m- to cool, to warm slightly, to take the chill off K.
-16. fül Av close fül-üm- to bring closer, to approximate, to bring near K. S.
-17. füy VI to squeeze (itself), to hold on (to itself) füy-üm- to tighten, to hold K.
-18. illku VI to get angry illku-m- to contradict, to make angry K.
-19. kofi VI to get heat up, to get overheat kofi-m- to heat, to overheat S.
-20. kon VI to enter, to begin kon-üm- to make come in, to add K. G.
-21. kudu VI to lie (down), to go to bed kudu-m- to lay out K.
-22. küme Aj good, nice, pleasant küme-m- to add value, to improve K.
-23. kümpu N piece, chunk, bit kümpu-m- to chop, to cut into pieces, to tear apart, to shred K.
-24. küpa VI to come küpa-m- to bring, to introduce K.
-25. la Aj dead, deceased lang-üm- to kill K. S. G.
-26. laf Aj flat, level lap-üm- to stretch, to extend S.
-27. law VI to lose hair/feathers law-üm- to shave, to shear, to pluck S.
-28. lef Aj fast (on going or moving) lep-üm- to make run (animals), to make flee, to make turn away S. G.
-29. (l)el VI to let oneself, to become loose, to detach (l)el-üm- to release, to let go, to kill K.
-30. lif Aj clean lip-üm- to clean K.
-31. llangkü VI to fall (down) llangkü-m- to drop, to lose K. S.
-32. lleg VI to grow (up), to come up llek-üm- to raise, to cultivate, to grow (plants), to bring out S. G.
-[…]
-34. llum VI to get hidden, to conceal oneself llum-üm- to hide, to conceal K.
-35. lüf Aj burned, roasted p-üm- to burn, to set fire, to roast S. G.
-36. lüg Aj white lüg-üm- to whiten, to bleach, to lighten (in color) S.
-37. lulu N noise, crash, upheaval lulu-m- to make noise, to cause crashing K.
-38. nag Av down, underneath nak-üm- to go down, to lower, to descend K. S. G.
-[…]
-40. neng VI to move (oneself) neng-üm- to move, to shake, to shudder, to direct K.
-41. neyküe VI to unleash, to get free neykü-m- to release, to let go, to free S.
-42. ngül VI to get gathered, to come together, to join ngül-üm- to gather, to bring together, to collect, to harvest K. S.
-43. nong VI to get broken down, to get damaged, to get scorched nong-üm- to spoil, to damage, to scorch K.
-44. nor Aj straight, correct, right nor-üm- to correct, to rectify, to straighten S.
-45. ñam VI to get lost, to get astray, to get loose ñam-üm- to lose, to release, to let go, to make disappear S. G.
-46. ñif VI to dry out ñip-üm- to dry S.
-47. ñom Aj meek, gentle, tame, docile, obedient ñom-üm- to tame, to domesticate, to tame K. S.
-48. pelof N light, clarity pelo-m- to illuminate, to light up, to turn on the light S.
-49. pewü VI to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m- to measure, to reach out using the arms, to twist, to turn K.
-50. piwü VI to dry out piwü-m- to dry (up) K. S.
-51. pod N dirt, filth, uncleanliness pod-üm- to dirty, to soil S.
-52. poñü N potato poñü-m- feed with potatoes K.
-53. püd Aj thick, dense, scattered, dispersed püd-üm- to spread, to scatter, to thicken, to extend, to spread out, to stretch, to ration, to distribute K. S.
-54. pümü Aj tense, tight, taut, strained pümü-m- to stretch, to tighten, to tense K.
-55. püna VI to get stuck, to get glued, to get adhered püna-m- to stick, to glue, to adhere S.
-56. püra VI to ascend, to go up, to climb püra-m-g to lift, to raise, to pick up, to keep in mind, to remember, to make evident, to highlight K. S. G.
-57. püsha VI to run out, to be finished püsha-m- to finish, to end, to spend, to use up, to exhaust G.
-58. pütre Aj burning, red-hot pütre-m- to smoke, to fumigateh K.
-59. puw VI to arrive, to reach, to achieve, to stay, to remain puw-üm- to send, to deliver, to convey, to serve K.
-60. rüngü VI to be pulverised, to be ground, to be crushed rüngü-m- to grind, to mill, to crush S.
-61. traf Aj together, joined, gathered, fitted trap-üm- to gather, to collect, to put together, to grab, to catch, to pick up K. S.
-62. tral VI to get hit, to get bumped, to get dented, to get bruised, to collide, to crash, to clash tral-üm- to hit, to strike, to beat K.
-63. tran Aj fallen (especially referring to trees) tran-üm- to make fall, to knock down, to knock over, to topple, to lay down G.
-64. trar VI to ooze, to suppurate, to secrete trar-üm- to drain the pus S.
-65. trem Aj grown, adult, mature trem-üm- to raise, to breed K. S. G.
-66. trof VI to burst, to break, to crack, to snap trop-üm- to split, to break S.
-67. trong Aj thick, dense trong-üm- to roof (houses)i K.
-68. trür Aj even, level, tight, composed trür-üm- to equalise, to compensate, to balance, to indemnify K. S.
-69. tüng VI to calm down, to take time tüng-üm- to leave alone, not to disturb K.
-70. üng Av while üng-üm- to wait K.
-[…]
-72. ürkü VI to get tired ürkü-m- to tire, to tire out G.
-[…]
-74. üy N name üy-üm- to name, to denominate S.
-75. wadkü VI to boil wadkü-m- to boil, to bring to a boil K. S. G.
-76. wallü VI to get surrounded, to get encompassed, to get completed wallü-m- to tell or recount in detail K.
-77. wef VI to appear, to show itself wep-üm- to make appear, to find, to show G.
-78. welli VI to get empty welli-m- to empty, to vacate K. S.
-79. wim Aj accustomed, used to, meek, tame, domesticated wim-üm- to make accustomed, to domesticate, to tame S.
-80. wüda VI to divide, to split up wüda-m- to divide, to distribute, to share, to split K. S.
-81. wünü VI to get stretched, to get elongated wünü-m- to stretch, to lengthen S.
-82. yall N son (of a man) yall-üm- to multiply, to engender S.
-83. yewe VI to feel embarrassed, to be ashamed yewe-m- to embarrass G.
-84. yung Aj pointed, sharp yung-üm- to sharpen, to make pointed K.
-a.
-In the ‘‘Collected from’’ column, the reference work containing examples with the corresponding root is identified.
-K.
-stands for Paskwal Koña in Wilhelm et al., 1930; S.
-for Smeets; and G.
-for Golluscio, 2007.
-b. The Mapuche use the term chaw-üm- ‘fathering’, or more literally ‘to cause to father’, to refer to ‘incubating eggs’.
-c.
-In reality, to convey the meaning of ‘to relate’, the compound che-m-ye- needs to be formed, which is analysed as -NN.che_person_people+CA.m34+TV.ye_to_carry.
+11. fa Dem this, here fa-m to be like this K. S.
+12. fay IV to ferment fay-üm to [make] ferment S.
+13. fe Dem that, there fe-m to be like that K. S.
+14. fey IV to get adapted, to fit fey-üm to adapt, to fit, to insert S.
+15. firkü Aj fresh, cool firkü-m to cool, to warm slightly, to take the chill off K.
+16. fül Av close fül-üm to bring closer, to approximate, to bring near K. S.
+17. füy IV to squeeze (itself), to hold on (to itself) füy-üm to tighten, to hold K.
+18. illku IV to get angry illku-m to contradict, to make angry K.
+19. kofi IV to get heat up, to get overheat kofi-m to heat, to overheat S.
+20. kon IV to enter, to begin kon-üm to make come in, to add K. G.
+21. kudu IV to lie (down), to go to bed kudu-m to lay out K.
+22. küme Aj good, nice, pleasant küme-m to add value, to improve K.
+23. kümpu N piece, chunk, bit kümpu-m to chop, to cut into pieces, to tear apart, to shred K.
+24. küpa IV to come küpa-m to bring, to introduce K.
+25. la Aj dead, deceased lang-üm to kill K. S. G.
+26. laf Aj flat, level lap-üm to stretch, to extend S.
+27. law IV to lose hair/feathers law-üm to shave, to shear, to pluck S.
+28. lef Aj fast (on going or moving) lep-üm to make run (animals), to make flee, to make turn away S. G.
+29. (l)el IV to let oneself, to become loose, to detach (l)el-üm to release, to let go, to kill K.
+30. lif Aj clean lip-üm to clean K.
+31. llangkü IV to fall (down) llangkü-m to drop, to lose K. S.
+32. lleg IV to grow (up), to come up llek-üm to raise, to cultivate, to grow (plants), to bring out S. G.
+33. lliw $\sim$ llüw $\sim$ lluw IV to get merged, to get melted, to get liquefied lliw-üm to merge, to melt, to liquefy G.
+34. llum IV to get hidden, to conceal oneself llum-üm to hide, to conceal K.
+35. lüf Aj burned, roasted p-üm to burn, to set fire, to roast S. G.
+36. lüg Aj white lüg-üm to whiten, to bleach, to lighten (in color) S.
+37. lulu N noise, crash, upheaval lulu-m to make noise, to cause crashing K.
+38. nag Av down, underneath nak-üm to go down, to lower, to descend K. S. G.
+39. nel Aj loose, untied, unleashed nel-üm $\sim$ nelk-üm to release, to let go, to free S.
+40. neng IV to move (oneself) neng-üm to move, to shake, to shudder, to direct K.
+41. neyküe IV to unleash, to get free neykü-m to release, to let go, to free S.
+42. ngül IV to get gathered, to come together, to join ngül-üm to gather, to bring together, to collect, to harvest K. S.
+43. nong IV to get broken down, to get damaged, to get scorched nong-üm to spoil, to damage, to scorch K.
+44. nor Aj straight, correct, right nor-üm to correct, to rectify, to straighten S.
+45. ñam IV to get lost, to get astray, to get loose ñam-üm to lose, to release, to let go, to make disappear S. G.
+46. ñif IV to dry out ñip-üm to dry S.
+47. ñom Aj meek, gentle, tame, docile, obedient ñom-üm to tame, to domesticate, to tame K. S.
+48. pelof N light, clarity pelo-m to illuminate, to light up, to turn on the light S.
+49. pewü IV to swirl, to gather in a whirl, to twist, to sprain, to coil, to curl up pewü-m to measure, to reach out using the arms, to twist, to turn K.
+50. piwü IV to dry out piwü-m to dry (up) K. S.
+51. pod N dirt, filth, uncleanliness pod-üm to dirty, to soil S.
+52. poñü N potato poñü-m feed with potatoes K.
+53. püd Aj thick, dense, scattered, dispersed püd-üm to spread, to scatter, to thicken, to extend, to spread out, to stretch, to ration, to distribute K. S.
+54. pümü Aj tense, tight, taut, strained pümü-m to stretch, to tighten, to tense K.
+55. püna IV to get stuck, to get glued, to get adhered püna-m to stick, to glue, to adhere S.
+56. püra IV to ascend, to go up, to climb püra-mg to lift, to raise, to pick up, to keep in mind, to remember, to make evident, to highlight K. S. G.
+57. püsha IV to run out, to be finished püsha-m to finish, to end, to spend, to use up, to exhaust G.
+58. pütre Aj burning, red-hot pütre-m to smoke, to fumigateh K.
+59. puw IV to arrive, to reach, to achieve, to stay, to remain puw-üm to send, to deliver, to convey, to serve K.
+60. rüngü IV to be pulverised, to be ground, to be crushed rüngü-m to grind, to mill, to crush S.
+61. traf Aj together, joined, gathered, fitted trap-üm to gather, to collect, to put together, to grab, to catch, to pick up K. S.
+62. tral IV to get hit, to get bumped, to get dented, to get bruised, to collide, to crash, to clash tral-üm to hit, to strike, to beat K.
+63. tran Aj fallen (especially referring to trees) tran-üm to make fall, to knock down, to knock over, to topple, to lay down G.
+64. trar IV to ooze, to suppurate, to secrete trar-üm to drain the pus S.
+65. trem Aj grown, adult, mature trem-üm to raise, to breed K. S. G.
+66. trof IV to burst, to break, to crack, to snap trop-üm to split, to break S.
+67. trong Aj thick, dense trong-üm to roof (houses)i K.
+68. trür Aj even, level, tight, composed trür-üm to equalise, to compensate, to balance, to indemnify K. S.
+69. tüng IV to calm down, to take time tüng-üm to leave alone, not to disturb K.
+70. üng Av while üng-üm to wait K.
+71. üre $\sim$ ürü IV to get wet üre-m $\sim$ ürü-m to wet, to moisten, to soak K. S.
+72. ürkü IV to get tired ürkü-m to tire, to tire out G.
+73. üy $\sim$ üyü IV to catch fire, to ignite üy-üm $\sim$ üyü-m to set fire, to ignite K. S.
+74. üy N name üy-üm to name, to denominate S.
+75. wadkü IV to boil wadkü-m to boil, to bring to a boil K. S. G.
+76. wallü IV to get surrounded, to get encompassed, to get completed wallü-m to tell or recount in detail K.
+77. wef IV to appear, to show itself wep-üm to make appear, to find, to show G.
+78. welli IV to get empty welli-m to empty, to vacate K. S.
+79. wim Aj accustomed, used to, meek, tame, domesticated wim-üm to make accustomed, to domesticate, to tame S.
+80. wüda IV to divide, to split up wüda-m to divide, to distribute, to share, to split K. S.
+81. wünü IV to get stretched, to get elongated wünü-m to stretch, to lengthen S.
+82. yall N son (of a man) yall-üm to multiply, to engender S.
+83. yewe IV to feel embarrassed, to be ashamed yewe-m to embarrass G.
+84. yung Aj pointed, sharp yung-üm to sharpen, to make pointed K.
+a The “Collected from” column indicates which source contains examples of the root with the causative suffix.
+K. = Wilhelm et al., 1930 (autobiography of longko Paskwal Koña); S. = Smeets, 2008 (A Grammar of Mapuche); G. = Golluscio, 2007 (Gramática descriptiva del mapuche).
+b The Mapuche use the term chaw-üm ‘fathering’, or more literally ‘to cause to father’, to refer to ‘incubating eggs’.
+c In reality, to convey the meaning of ‘to relate’, the compound che-m-ye needs to be formed, which is analysed as -nn.che_person_people+ca.m+tv.ye_to-carry.
 This is the way to express a relationship with someone in Mapudüngun.
 For example, laku-ye-fi-ñ translates to ‘he is my paternal grandfather’, which literally means ‘I carry him as a grandfather’ (Becerra, 2011: 115).
-d.
-In the case of interrogatives like ‘‘what, which, how’’, etc., staring with chu-, these are formed by adding other suffixes.
+d In the case of interrogatives like “what, which, how”, etc., starting with chu-, these are formed by adding other suffixes.
 For example, chu-m-al means ‘what for’, chu-m-nge-chi means ‘how’, chu-m-ül means ‘when’, and so on.
-e.
-It is very likely that neykü is a lexicalised form, composed of the root nel- listed in line 39, and the unproductive suffix -kü-, which is found in a few verbal roots but has not been documented or described in existing literature.
-f. pelo is a lexicalised form of a nominalised verb (see example E.28.a in Chandía, 2025).
-g.
-The form püra-m- generates compounds with many action verbs to indicate an upward direction, such as in tralkatunpüramün, which means ‘to shoot into the air’.
+e It is very likely that neykü is a lexicalised form, composed of the root nel- listed in line 39, and the unproductive suffix -kü, which is found in a few verbal roots but has not been documented or described in existing literature.
+f pelo is a lexicalised nominalised verb, likely from pe- ‘sight, vision’ + fossilised suffix -lo (no longer productive), meaning ‘light, clarity’.
+For the origin of similar lexicalised forms, see footnote 18.
+For an example of pelo in a verbal form, see the companion valency article (Chandía, 2026, E.28a).
+g The form püra-m generates compounds with many action verbs to indicate an upward direction, such as in tralkatunpüramün, which means ‘to shoot into the air’.
 It is also used figuratively, as in kimpüramün, meaning ‘to become aware’; literally: ‘to raise knowledge’.
-h. When pütre is nominalised as pütrem -AJ.pütre_blazing+CA.m34+NOM.ø, it means ‘cigar’ or ‘tobacco’.
-i. trongum- is to thatch a ruka (a traditional Mapuche house) with reed or tall grass, these materials must be placed very densely, ensuring a good level of thickness.
+h When pütre is nominalised as pütrem (-aj.pütre_blazing+ca.m+nom.ø), it means ‘cigar’ or ‘tobacco’.
+i trongum- is to thatch a ruka (a traditional Mapuche house) with reed or tall grass; these materials must be placed very densely, ensuring a good level of thickness.
+?refname?
+  • A.
+Álvarez, R.
+Aranovic, R.
+Brown, J.
+Carbonell, C.
+Fasola, A.
+Lavie, L.
+Levin, A.
+Font-Llitjos, C.
+Monson, E.
+Peterson, K.
+Probst, and R.
+Vega (2005) Informe final del Proyecto Avenue / Mapudungun: Desarrollo de herramientas informáticas para el mapudüngun que se habla en Chile.
+Technical report Language Technologies Insitute.
+Carnegie Mellon University, Pensilvania, EE.
+UU..
+Cited by: §1.1.2.
+  • R.
+Becerra (2011) Una propuesta funcional y tipológica de descriptores sintácticos para el mapudungun.
+Alpha (Osorno), pp.
+111–125 (es).
+Note: PDF External Links: ISSN 0718-2201, Link Cited by: §4.2.
+  • A.
+Chandía (2012) Düngupeyem_alfa_v0.1: un prototipo de analizador morfológico para el mapudüngun a través de transductores de estados finitos.
+Máster en Ciencia Cognitiva y Lingüística, Universidad de Barcelona.
+Note: PDF Cited by: §1.
+  • A.
+Chandía (2014) CORLEXIM Corpus lexicográfico del mapudüngun.
+Note: Online resource http://www.chandia.net/corleximAvailable dictionaries: 1.
+Valdivia, L.
+de (1606): mapudüngun-castellano.
+2.
+Febrés, A.
+(1765): castellano-mapudüngun.
+3.
+Febrés, A.
+([1765] 1882): mapudüngun-castellano.
+4.
+Febrés, A.
+(1846a): mapudüngun-castellano.
+5.
+Febrés, A.
+(1846b): castellano-mapudüngun.
+6.
+Augusta, F.
+de (1916b): castellano-mapudüngun.
+7.
+Augusta, F.
+de (1916a): mapudüngun-castellano.
+8.
+Smeets, I.
+(2008): mapudüngun-inglés Cited by: item 1, item 2, item 3, item 4, §1.1.3, §1.1.3, §1.1.3, §1.1.3, §1.1.3, §2.1, footnote 5.
+  • A. Chandía (2021) FST Morphological Analyser and Generator for Mapudüngun. arXiv. Note: PDF External Links: Document, 2109.09176 Cited by: §1, §1, footnote 4.
+  • A.
+Chandía (2022) A Mapudüngun FST Morphological Analyser and its Web Interface.
+In Proceedings of the Language Resources and Evaluation Conference, Marseille, France, pp.
+6540–6547.
+Note: PDF Cited by: §1.
+  • A.
+Chandía (2026) Valency classification of mapudungun verbal roots.
+established by the language’s own morphotactics.
+External Links: 2604.00789, Link Cited by: §1.2, §1.2, §1, §1, item 6., §2.2, §2.3, §2.5, §2, §3, §3, §4.2, §4.2.
+  • F. de Augusta (1903) Gramática araucana. Imprenta Central, J. Lampert, Valdivia. Note: PDF Cited by: item d..
+  • L.
+Golluscio (2007) Morphological causatives and split intransitivity in mapudungun.
+International Journal of American Linguistics 73 (2), pp.
+209–238.
+External Links: ISSN 00207071, 15457001, Link Cited by: §2.5, §2.5, §4.2.
+  • E. Longkon (2011) Morfología y aspectos del mapudüngun. Biblioteca de signos, Universidad Autónoma Metropolitana, México. External Links: ISBN 9786074775532 Cited by: §2.5.
+  • C.
+Martinez and M.
+Carmona (2023) El Vocabulario de Bernard Havestad (1777 – Pars Quarta), versión mapudungun-español.
+.
+Note: Online resource https://benmolineaux.github.io/bookshelf/ Cited by: footnote 6.
+  • I.
+Smeets (2008) A Grammar of Mapuche.
+De Gruyter Mouton, Berlin, Boston.
+Note: PDF External Links: Document, ISBN 9783110211795 Cited by: §1, item a., item b., item c., 1st item, §1.2, §1, §1, §1, §1, §1, §2.1, §2.5, §2.5, §3, §4.2, footnote 2.
+  • B.
+Villena (2016) Innovación léxica en mapudungún: genuinidad, productividad y planificación.
+Doctoral Dissertation, Universitat Pompeu Fabra.
+External Links: Link Cited by: §1.1.3, §1.1.3, §1.1.3, §1.1.3, footnote 12.
+  • A.
+Wilhelm, Pascual.
+Koña, and R.
+Lenz (1930) Vida y costumbres de los indígenas araucanos en la segunda mitad del siglo XIX.
+Imprenta Universitaria, Santiago, Chile.
+Note: PDF External Links: LCCN 37025221 Cited by: 2nd item, §1.2, §2.1, §2.3, §2.5, §3, §4.2.
+  • F. Zúñiga (2006) Mapudüngun. el habla mapuche. Centro de Estudios Públicos, Santiago, Chile. Note: PDF External Links: ISBN 9783895869761 Cited by: footnote 10.